### v2

# FIT-GNN: Faster Inference Time for GNNs Using Coarsening

Abstract
Scalability of Graph Neural Networks (GNNs) remains a significant challenge, particularly when dealing with large-scale graphs. To tackle this, coarsening-based methods are used to reduce the graph into a smaller graph, resulting in faster computation. Nonetheless, prior research has not adequately addressed the computational costs during the inference phase. This paper presents a novel approach to improve the scalability of GNNs by reducing computational burden during both training and inference phases. We demonstrate two different methods (Extra-Nodes and Cluster-Nodes). Our study also proposes a unique application of the coarsening algorithm for graph-level tasks, including graph classification and graph regression, which have not yet been explored. We conduct extensive experiments on multiple benchmark datasets in the order of $100K$ nodes to evaluate the performance of our approach. The results demonstrate that our method achieves competitive performance in tasks involving classification and regression on nodes and graphs, compared to traditional GNNs, while having single-node inference times that are orders of magnitude faster. Furthermore, our approach significantly reduces memory consumption, allowing training and inference on low-resource devices where traditional methods struggle.
$\dagger$\$$\dagger$\$footnotetext: These authors contributed equally to this work
[…]
images/img1
Figure 1: The figure shows how the coarsening algorithm reduces a graph $G$ with $15$ nodes to a coarsened graph $G^{\prime}$ with $3$ nodes. It also shows which nodes belong to which partition.
1 Introduction
Graph Neural Networks (GNNs) have proven to be a tool of surprising versatility and modeling capacity. However, there are still a few issues that limit their general applicability, the chief among these being the scalability of the GNN models. This scalability issue is problematic for both the training and the inference phases. Previous research has attempted to address this by reducing the graph into a smaller graph in order to reduce the computational cost. One such broad technique is graph coarsening methods, which judiciously remove nodes and edges from the graph. However, there are still certain points of weakness of these algorithms. For one, coarsening methods can potentially remove important node information. Secondly, and more pertinent to the current discussion, coarsening methods only reduce the training computation; they do not address the inference time. The next obvious question would be to ask why we cannot use coarsening methods for inference as well. Given a graph $G$, coarsening algorithms create a coarsened graph $G^{\prime}$. While $G^{\prime}$ is smaller than $G$, the nodes in $G^{\prime}$ are combinations of nodes in $G$, which can be both train and test nodes. As a result, the prediction per test node is not available. In another scenario, suppose a node $v^{\prime}\in G^{\prime}$ is a combination of $v_{i},v_{j},v_{k}\in G$ where $v_{i},v_{j},v_{k}$ are test nodes. Also, let us assume that the classes of $v_{i},v_{j},v_{k}$ are $0,0,2$ respectively. The coarsening algorithm will take the majority label and assign class 00 to $v^{\prime}$. Hence, the model will be trained to predict only the class 00, leading to discarding the quantification of the model’s performance on predicting less represented nodes. This is the research gap that we seek to address in this work. We use coarsening algorithms to partition the graph into a set of subgraphs $\mathcal{G}_{s}=\{G_{1},G_{2},\dots,G_{k}\}$. We show methods to train on these subgraphs that show equivalent results to non-coarsened baselines and sometimes perform better. This comes hand in hand with a significant reduction in inference time and memory consumption.
We demonstrate two different methods to append additional nodes in the subgraphs— Extra Nodes and Cluster Nodes that respectively append additional nodes to each subgraph, either by adding neighboring nodes, or by adding representational nodes corresponding a neighboring cluster. We theoretically characterize the upper bound of the coarsening ratio for which the time complexity of our approach is better than the naive uncoarsened approach. In node classification and node regression tasks, the baselines typically require the entire dataset during inference to get a prediction for a single node. In contrast, our method only needs to pass the subgraph to which the node belongs. Our results show up to $\mathbf{100\times}$ reduction in inference time and up to $\mathbf{70\times}$ less memory than the baseline while achieving equivalent or better performance.
2 Preliminaries
2.1 Graph Notations
Given an undirected graph $G=(V,E,X,W)$, $V$ is the vertex set, $E$ is the edge set, $X\in\mathbb{R}^{n\times d}$ is the feature matrix, and $W$ is the edge weight matrix. The feature of each node $v_{i}$ in the graph is the $i$th row vector in $X$ represented as $x_{i}\in\mathbb{R}^{d}$. Let $|V|=n$ be the number of nodes and $|E|=m$ be the number of edges. Let us assume $A\in\{0,1\}^{n\times n}$ be the adjacency matrix of $G$, $d_{i}$ to be the degree of the node $v_{i}$. $D\in\mathbb{R}^{n\times n}$ is a diagonal degree matrix with $i$th diagonal element as $d_{i}$. We denote $\mathcal{N}_{j}(v_{i})$ as the $j$-hop neighbourhood of node $v_{i}\in V$.
2.2 Graph Neural Network
Graph Neural Network is a neural network designed to work for graph-structured data. The representation of each node in the graph is updated recursively by aggregation and transforming node representations of its neighbors. Kipf and Welling (2017) introduced Graph Convolutional Network (GCN) as follows:
$X_{l+1}=\sigma(\tilde{D}^{-\frac{1}{2}}\tilde{A}\tilde{D}^{-\frac{1}{2}}X_{l}\mathcal{W}_{l})$ (1)
where $X_{l}$ is the node representation after $l$ layers of GCN, $\tilde{A}=A+I$, $\tilde{D}=D+I$, $I$ is an identity matrix of same dimension as $A$. $\mathcal{W}_{l}$ is the learnable parameter and $\sigma$ is a non-linear activation function.
2.3 Graph Coarsening and Graph Partitioning
Loukas (2018) discussed multiple coarsening algorithms which create a coarsened graph $G^{\prime}=(V^{\prime},E^{\prime},X^{\prime},W^{\prime})$ from a given graph $G=(V,E,X,W)$. We refer the vertex set $V^{\prime}=\{v^{\prime}_{1},v^{\prime}_{2},\dots,v^{\prime}_{k}\}$ of $G^{\prime}$ as coarsened nodes. Given a coarsening ratio $r\in[0,1]$, we have $k=n\times r$. We can interpret it as creating $k$ disjoint partitions, $C_{1},C_{2},\dots,C_{k}$, from a graph of $n$ nodes. Mathematically, we create a partition matrix $P\in\{0,1\}^{n\times k}$, where $P_{ij}=1$ if and only if node $v_{i}\in C_{j}$.
For the coarsened graph, the adjacency matrix $A^{\prime}=P^{T}AP$. Similarly, the corresponding degree matrix $D^{\prime}=P^{T}DP$. Huang et al. (2021) used normalized partition matrix $\mathcal{P}=PC^{-\frac{1}{2}}$, where $C$ is defined as a diagonal matrix with diagonal entries $C_{j,j}=|C_{j}|,j=1,2,\dots,k$.
Along with a coarsened graph, we create a set of disjoint subgraphs $\mathcal{G}_{s}=\{G_{1},G_{2},\dots,G_{k}\}$ corresponding to the partitions $C_{1},C_{2},\dots,C_{k}$. The number of nodes in each partition $C_{i}$ is denoted by $n_{i}$. Each subgraph $G_{i}$ is the induced subgraph of $G$ formed by the nodes in $C_{i}$. $A_{i}$ and $D_{i}$ are the adjacency matrix and degree matrix of $G_{i}$ respectively.
3 Related Work
The goal of reducing computation costs has been tackled via different methodologies. Research like Chiang et al. (2019) and Zeng et al. (2020) use a subgraph sampling-based approach. They sample a small subgraph at each training iteration and train the GNN. Other techniques like layer-wise sampling along with mini-batch training have been studied by  Chen et al. (2018); Chen and Zhu (2018); Cong et al. (2020); Zou et al. (2019). Fahrbach et al. (2020) introduced a new coarsening algorithm based on Schur complements to generate embeddings for vertices in a large graph. Huang et al. (2021) introduced using GNN for a coarsened graph. The motivation was to utilize graph coarsening algorithms to reduce the number of nodes and edges in the graph, resulting in less training time. This approach helps to train GNN for graphs with a large number of nodes. However, inference time is not reduced because the nodes cannot be removed during inference. This training process also caused the removal of important node-level information. Later, Xue et al. (2022) explored shifting the training process from graph-level training to subgraph-level. Coarsening algorithms were used to split the graph into $k-$subgraphs, resulting in the removal of a few edges. Further, to speed up the training process, they propose a multi-device training approach where subgraphs are sent to different devices. This results in a faster training process. Kumar et al. (2023) further proposed the usage of feature-based coarsening algorithms.
4 Methodology
In Huang et al. (2021), a graph $G=(V,E,X,W)$ is reduced to $G^{\prime}=(V^{\prime},E^{\prime},X^{\prime},W^{\prime})$ using a partition matrix $P$ generated from the coarsening algorithm. Here $X^{\prime}=P^{T}X$. The labels for the coarsened graph are $Y^{\prime}=\arg\max(P^{T}Y)$, where $Y$ is the label matrix of $G$ storing the one-hot encoding of the label of each node for the node classification task. While using $\arg\max$, some essential information is lost. Therefore, we follow a subgraph-level training process where we don’t remove these label information. To explore another node-level task, such as node regression, functions such as the mean of regression targets wouldn’t be appropriate because they lose the variance of the targets. Hence, our goal is to use subgraphs for training and inference purposes.
To tackle the issue of loss of information corresponding to the nodes (and their edges) present at the periphery of the subgraphs after partitioning the graph, we append nodes in two ways:
[…]
images/img2
Figure 2: This figure shows the methods of appending additional nodes in $G_{1},G_{2},G_{3}$.
$\bullet$ Extra Nodes: Adding the $1$-hop neighbouring nodes in each subgraph $G_{i}$ denoted by $\mathcal{E}_{G_{i}}$.
$\mathcal{E}_{G_{i}}=\bigcup_{v\in G_{i}}\left\{u\in V:u\in\mathcal{N}_{1}(v)\wedge u\notin G_{i}\right\}$
However, we don’t perform backpropagation of the prediction on $\mathcal{E}_{G_{i}}$. Xue et al. (2022) introduced this approach to reduce information loss after partitioning. An edge between two Extra Nodes in $\mathcal{E}_{G_{i}}$ is added if they were connected in $G$.
Algorithm 1 Node-Model$(L,A,D,X)$
1:Number of Layers $L$; $A$; $D$; $X_{0}=X$;
2:for $i=1$ to $L$ do
[…]
4:end for
5:$Z=X_{L}\mathcal{W}_{L}$
6:return $Z$
$\bullet$ Cluster Nodes: Although adding $1$-hop neighboring nodes in the subgraph helps recover lost information after partition, for GNN models with a large number of layers, information loss will still persist. Instead of adding the neighboring nodes, we will add a node, namely Cluster Nodes, representing the neighboring cluster, denoted by $\mathcal{C}_{G_{i}}$.
$\mathcal{C}_{G_{i}}=\bigcup_{v\in\mathcal{E}_{G_{i}}}\left\{t:P_{v,t}\neq 0\right\}$
Liu et al. (2024) introduced this method to add nodes in the graph instead of adding to separate subgraphs. They also added cross-cluster edges among these Cluster Nodes and proposed a subgraph sampling-based method after adding these Cluster Nodes and cross-cluster edges. In our work, we add cross-cluster edges; however, we do not sample from the constructed set of subgraphs.
Figure 2 illustrates appending these additional nodes using these approaches. The dashed line shows which nodes are added as part of these additional nodes. Predictions on these newly appended nodes will not contribute to loss calculation and weight update.
Lemma 1.
Models with $1$ layer of GNN cannot distinguish between $G$ and $\mathcal{G}_{s}$ when Extra Nodes method is used.
Proof.
There are $2$ sets of nodes in $G_{i}$ as follows:
  • The set $S_{1}$ of nodes with 1-hop neighbours in $G_{i}$.
  • The set $S_{2}$ of nodes with not all 1-hop neighbors in $G_{i}$
Let $\mathcal{I}^{1}_{i}$ be the number of nodes whose information does not get passed on after $1$ layer of GNN for $G_{i}$.
$\mathcal{I}^{1}_{i}=\left|\bigcup_{v\in S_{2}}\mathcal{N}_{1}(v)-V(G_{i})\right|$
Now, |$\begin{split}\textup{Now, }\left|\bigcup_{v\in S_{1}}\left\{u\in V:u\in\mathcal{N}_{1}(v)\wedge u\notin G_{i}\right\}\right|=0\\ \textup{Also, }\mathcal{E}_{G_{i}}=\bigcup_{v\in S_{2}}\left\{u\in V:u\in\mathcal{N}_{1}(v)\wedge u\notin G_{i}\right\}\\ \end{split}$
$\Rightarrow\left|\mathcal{E}_{G_{i}}\right|=\mathcal{I}^{1}_{i}$
Hence when Extra Nodes is used, 1 Layer GNN model cannot distinguish $G$ and $\mathcal{G}_{s}$. ∎
Algorithm 2 Training GNN (Train on $\mathcal{G}_{s}$)
1:$G=(V,E,X)$; Labels $Y$; Model $M$; Loss $\ell$;
2:Apply a coarsening algorithm on $G$, and output a normalized partition matrix $P$.
3:Construct $\mathcal{G}_{s}=\{G_{1},G_{2},\dots,G_{k}\}$ using P;
4:Construct mask for each subgraph $Mask=\{mask_{1},mask_{2},\dots,mask_{k}\}$. Masks denote the nodes that are used for training;
5:Construct feature matrix $\{X_{1},X_{2},\dots,X_{k}\}$ for $\mathcal{G}_{s}$;
6:Construct label matrix $\{Y_{1},Y_{2},\dots,Y_{k}\}$ for $\mathcal{G}_{s}$ ;
7:$O\leftarrow[]$; $Y\leftarrow[]$;
8:for each subgraph $G_{i}$ in $\mathcal{G}_{s}$ do
9:     $O_{i}=M(A_{i},D_{i},X_{i})$;
10:     $O.append(O_{i}[mask_{i}])$;
11:     $Y.append(Y_{i}[mask_{i}])$;
12:end for
13:$Loss=\ell(O,Y)$;
14:Train $M$ to minimize $Loss$;
Using Cluster Nodes approach to append additional nodes to subgraphs is better for two reasons:
$\bullet$ The number of nodes added to a subgraph via Cluster Nodes is less than or equal to the number of nodes added via the Extra Nodes approach. This is because, using Extra Nodes, we add all the neighboring nodes. However, we only add one representative node for each neighboring cluster using Cluster Nodes. As the GNN propagation depends on the number of nodes, the time taken to train and infer these modified subgraphs with Cluster Nodes is less than or equal to that with Extra Nodes.
$\bullet$ As mentioned before, the Extra Nodes approach reduces the information loss due to partitioning. However, with multiple layers of GNN, longer dependencies are not captured. In contrast, the Cluster Nodes approach overcomes this by computing a single Cluster Nodes’s representation as a combination of features of all the nodes present in the corresponding cluster. This results in sharing further node information in $1$-hop. Additionally, the transfer of information from one cluster to another is captured.
4.1 Node-level Task
We use Algorithm 1 with $L$ number of layers to create a standard GNN model for node-level tasks. Training the model on $G^{\prime}$ is similar to the Huang et al. (2021) training algorithm. We design Algorithm 2 to train the model on $\mathcal{G}_{s}$.
For the node classification task, $\arg\max$ is used for creating labels for $G^{\prime}$ and CrossEntropy as a loss function. For the node regression task, we do not create any coarsened graph. Mean Absolute Error (MAE) is used as the loss function.
4.2 Graph-level Task
Unlike node-level tasks, the number of dataset samples is not decreasing. For a given graph $G\in\mathcal{D}$, we create $G^{\prime}$ and $\mathcal{G}_{s}=\{G_{1},G_{2},\dots,G_{k}\}$. Each subgraph contains overlapping nodes when we use Extra Nodes or Cluster Nodes. Therefore, a boolean array $mask_{i}$ is created for each subgraph $G_{i}$ such that $True$ is set for nodes that actually belong to the subgraph $G_{i}$ (not as an Extra Nodes or Cluster Nodes); otherwise, $False$.
Unlike node-level tasks, we cannot train and infer on $G^{\prime}$ and $\mathcal{G}_{s}$ using the same model. We use Algorithm 3 to create a model to train on $G^{\prime}$. We design Algorithm 4 to train the model on $\mathcal{G}_{s}$. We use the CrossEntropy loss function for the graph classification task and the Mean Absolute Error(MAE) for the graph regression task. We use Algorithm 2 without the backpropagation for inference.
Algorithm 3 Graph-Model-$G^{\prime}(L,A^{\prime},D^{\prime},X^{\prime})$
1:Number of Layers $L;A^{\prime};D^{\prime};X_{0}=X^{\prime}$
2:for $i=1$ to $L$ do
[…]
4:end for
5:$\bar{X}=\textup{MaxPooling}(X_{L})$
6:$Z=\bar{X}\mathcal{W}_{L}$
7:return $Z$
Algorithm 4 Graph-Model-$\mathcal{G}_{s}(\mathcal{G}_{s},L,mask)$
1:$\mathcal{G}_{s}=\{G_{1},G_{2},\dots,G_{k}\}$; Number of Layers $L$; $mask=\{mask_{1},mask_{2},\dots,mask_{k}\}$;
2:$X\leftarrow[]$
3:for each subgraph $G_{i}\in\mathcal{G}_{s}$ do
4:     for $j=1$ to $L$ do
5:         $X_{j}=\sigma(\tilde{D_{i}}^{-\frac{1}{2}}\tilde{A_{i}}\tilde{D_{i}}^{-\frac{1}{2}}X_{j-1}\mathcal{W}_{j-1})$
6:     end for
7:     $X.append(X_{L}[mask_{i}])$
8:end for
9:$\bar{X}=\textup{MaxPooling}(X)$
10:$Z=\bar{X}\mathcal{W}_{L}$
11:return $Z$
Train Test
Classical $nd^{2}+n^{2}d$ $nd^{2}+n^{2}d$
Coarsening $kd^{2}+k^{2}d$ $nd^{2}+n^{2}d$
FIT-GNN $\begin{matrix}kd^{2}+k^{2}d\\ +\sum_{i=1}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]\end{matrix}$ $\begin{matrix}\sum_{i=1}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]\end{matrix}$
Table 1: This table shows the training and inference time complexity of FIT-GNN as compared to classical and coarsening approaches.
Train Test
Classical $n^{2}+nd+d^{2}$ $n^{2}+nd+d^{2}$
Coarsening $k^{2}+kd+d^{2}$ $n^{2}+nd+d^{2}$
FIT-GNN $\begin{matrix}k^{2}+kd+d^{2}\\ +\max_{i=1}[\bar{n}_{i}^{2}+\bar{n}_{i}d]\end{matrix}$ $\begin{matrix}d^{2}\\ +\max_{i=1}[\bar{n}_{i}^{2}+\bar{n}_{i}d]\end{matrix}$
Table 2: This table shows the training and inference space complexity of FIT-GNN compared to classical and coarsening approaches.
4.3 Time and Space Complexity
The dimensions of the matrices in Equation 1 are $(n\times n)$ for $\tilde{A}$, $(n\times d)$ for $X_{l}$ and $(d\times d)$ for $\mathcal{W}_{l}$. Time complexity for one layer of GNN computation is $\mathcal{O}(n^{2}d+nd^{2})$. If we take $L$ layers, then the total time is $\mathcal{O}(Ln^{2}d+Lnd^{2})$. The space complexity is $\mathcal{O}(n^{2}+Lnd+Ld^{2})$. When we compute on a sparse graph, the time complexity is $\mathcal{O}(m+Lnd^{2})$ and space complexity is $\mathcal{O}(m+Lnd+Ld^{2})$.
Huang et al. (2021) improved the time and space complexity (mentioned in Table 1 and Table 2) for training the network by reducing the number of nodes from $n$ to $k$. However, the inference time and space complexity remain the same.
Let us compare three different models. One is the classical model, where no coarsening or partitioning is done; the second is the Huang et al. (2021) approach; third is our approach where from the graph $G$ we create $G^{\prime}$ and $\mathcal{G}_{s}$.
The inference time complexity of our model is $\sum_{i=1}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]$, where $\bar{n}_{i}=n_{i}+\phi_{\textup{max}}$ where $\phi_{\textup{max}}=\max_{i}(\mathcal{E}_{G_{i}})$ for Extra Nodes or $\phi_{\textup{max}}=\max_{i}(\mathcal{C}_{G_{i}})$ for Cluster Nodes is the maximum number of additional nodes added per subgraph. Table 1 and Table 2 show the comparison between different approaches.
Lemma 2.
The inference time complexity is at most $\sum_{i=1}^{k}[(n_{i}+\phi_{\textup{max}})^{2}d+(n_{i}+\phi_{\textup{max}})d^{2}]\leq n^{2}d+nd^{2}$ for $r\leq\frac{d-2}{d+\phi_{\textup{max}}}$ and $\phi_{\textup{max}}\leq\frac{n^{2}-\sum_{i=1}^{k}n_{i}^{2}}{nd}$
Proof.
$\sum_{i=1}^{k}[(n_{i}+\phi_{\textup{max}})^{2}d+(n_{i}+\phi_{\textup{max}})d^{2}]\\ =\sum_{i=1}^{k}[n_{i}^{2}d+\phi_{\textup{max}}^{2}d+2n_{i}\phi_{\textup{max}}d+n_{i}d^{2}+\phi_{\textup{max}}d^{2}]\\ =k\phi_{\textup{max}}^{2}d+2n\phi_{\textup{max}}d+nd^{2}+k\phi_{\textup{max}}d^{2}+\sum_{i=1}^{k}n_{i}^{2}d\\ =nr\phi_{\textup{max}}^{2}d+2n\phi_{\textup{max}}d+nd^{2}+nr\phi_{\textup{max}}d^{2}+\sum_{i=1}^{k}n_{i}^{2}d\\ (\textup{where }k=nr)\\ =nd^{2}+n\phi_{\textup{max}}d(r\phi_{\textup{max}}+2+rd)+\sum_{i=1}^{k}n_{i}^{2}d$
Now we know, $r\leq\frac{d-2}{d+\phi_{\textup{max}}}\Rightarrow r\phi_{\textup{max}}+2+rd\leq d$ and $\phi_{\textup{max}}\leq\frac{n^{2}-\sum_{i=1}^{k}n_{i}^{2}}{nd}\Rightarrow n\phi_{\textup{max}}d\leq n^{2}-\sum_{i=1}^{k}n_{i}^{2}$.
Using this, we can say,
$\begin{split}&nd^{2}+n\phi_{\textup{max}}d(r\phi_{\textup{max}}+2+rd)+\sum_{i=1}^{k}n_{i}^{2}d\\ \leq&nd^{2}+(n^{2}-\sum_{i=1}^{k}n_{i}^{2})d+\sum_{i=1}^{k}n_{i}^{2}d\\ =&nd^{2}+n^{2}d\end{split}$
Hence, $\sum_{i=1}^{k}[(n_{i}+\phi_{\textup{max}})^{2}d+(n_{i}+\phi_{\textup{max}})d^{2}]\leq n^{2}d+nd^{2}$. ∎
For a given graph, if the above condition is satisfied, our approach has better inference time and space complexity than other approaches.
5 Experiments
Once we have constructed $G^{\prime}$ and $\mathcal{G}_{s}$, we train FIT-GNN in 4 different setups:
$\bullet$ Gc-train-to-Gs-train: Train the GNN model on $G^{\prime}$, then use the learned weight as an initialization for training and final inference on $\mathcal{G}_{s}$.
$\bullet$ Gc-train-to-Gs-infer: Train the GNN model on $G^{\prime}$, then infer on $\mathcal{G}_{s}$.
$\bullet$ Gs-train-to-Gs-infer: Train and infer only on $\mathcal{G}_{s}$.
$\bullet$ Gc-train-to-Gc-infer: Unlike node-level tasks, we can use $G^{\prime}$ to infer for graph-level tasks. Therefore, in this setup, we train and infer on $G^{\prime}$.
For tasks like node regression, we can only perform Gs-train-to-Gs-infer because a coarsened graph is not created.
5.1 Experimental Setup
Dataset descriptions are in Section A of the appendix, and Section C of the appendix covers the hyperparameters, device configuration, and key packages used for the experiments. Source Code present at https://github.com/Roy-Shubhajit/FIT-GNN.
5.1.1 Node regression
We use three real-world network datasets for the node regression task: Chameleon, Crocodile, and Squirrel Rozemberczki et al. (2019). The nodes are divided into train (30%), validation (20%), and test (50%) sets. We perform 20 runs and report the average and standard deviation of the smallest 10 normalized MAE, with normalization achieved by dividing the loss by the standard deviation of the target values.
5.1.2 Node classification
We conducted node classification experiments on six real-world network datasets: Cora McCallum et al. (2000), Citeseer Giles et al. (1998), Pubmed Sen et al. (2008), Coauthor Physics Shchur et al. (2018), DBLP Tang et al. (2008), and a subset of OGBN-Products Hu et al. (2020) extracted using Leiden algorithm Traag et al. (2019). We used three types of splits for the datasets: we applied the public split from Yang et al. (2016) for Cora, Citeseer, and PubMed. In the “few” split, we set aside labeled nodes per class for training and validation, while the rest form the test set. We ran 20 trials for each configuration and reported the average and standard deviation of the top 10 accuracies.
r Dataset
Chameleon Crocodile Squirrel
Classical - $0.844\pm 0.000$ $0.853\pm 0.000$ $0.809\pm 0.000$
FIT-GNN $0.1$ $\mathbf{0.460\pm 0.002}$ $\mathbf{0.362\pm 0.002}$ $\mathbf{0.648\pm 0.002}$
$0.3$ $0.499\pm 0.006$ $0.364\pm 0.001$ $0.672\pm 0.003$
Table 3: This table shows the normalized MAE loss (lower is better) for node regression tasks on different datasets with Gs-train-to-Gs-infer experiment setup and Cluster Nodes method.
Fixed Few
r Huang et al. (2021) FIT-GNN Huang et al. (2021) FIT-GNN
$0.1$ $0.772\pm 0.006$ $\mathbf{0.829\pm 0.005}$ $0.676\pm 0.051$ $\mathbf{0.695\pm 0.016}$
$0.3$ $\mathbf{0.817\pm 0.016}$ $0.815\pm 0.005$ $0.694\pm 0.045$ $\mathbf{0.719\pm 0.011}$
$0.5$ $\mathbf{0.827\pm 0.001}$ $0.807\pm 0.005$ $0.688\pm 0.045$ $\mathbf{0.729\pm 0.007}$
$0.7$ $\mathbf{0.824\pm 0.002}$ $0.800\pm 0.008$ $0.679\pm 0.046$ $\mathbf{0.706\pm 0.012}$
Table 4: Results on Cora Dataset. The metrics shown here include the average and standard deviation of the top 10 accuracies over 20 runs. We show Gc-train-to-Gs-train experimental setup result for fixed and Gc-train-to-Gs-infer experimental result for few. The highest accuracy for each coarsening ratio is shown in bold.
$\mathbf{r=0.1}$ $\mathbf{r=0.3}$
Dataset Split Huang et al. (2021) FIT-GNN Huang et al. (2021) FIT-GNN
Citeseer Few $0.583\pm 0.063$ $\mathbf{0.614\pm 0.021}$ $0.581\pm 0.052$ $\mathbf{0.624\pm 0.022}$
Fixed $\mathbf{0.711\pm 0.004}$ $0.677\pm 0.023$ $\mathbf{0.714\pm 0.003}$ $0.703\pm 0.007$
Pubmed Few $\mathbf{0.685\pm 0.052}$ $0.627\pm 0.116$ $\mathbf{0.687\pm 0.042}$ $0.676\pm 0.038$
Fixed $\mathbf{0.783\pm 0.005}$ $0.704\pm 0.022$ $\mathbf{0.784\pm 0.004}$ $0.677\pm 0.013$
DBLP Few $0.679\pm 0.056$ $\mathbf{0.683\pm 0.072}$ $0.648\pm 0.052$ $\mathbf{0.670\pm 0.034}$
Random $0.760\pm 0.021$ $\mathbf{0.789\pm 0.005}$ $0.745\pm 0.019$ $\mathbf{0.771\pm 0.007}$
Physics Coauthor Few $\mathbf{0.878\pm 0.036}$ $0.783\pm 0.028$ $\mathbf{0.908\pm 0.023}$ $0.877\pm 0.019$
Random $\mathbf{0.915\pm 0.014}$ $0.830\pm 0.048$ $\mathbf{0.934\pm 0.006}$ $0.910\pm 0.011$
OGBN-Products Random $0.296\pm 0.034$ $\mathbf{0.346\pm 0.014}$ $0.316\pm 0.016$ $\mathbf{0.406\pm 0.009}$
Table 5: Results for node classification tasks with accuracy as the metric (higher is better). We use the Cluster Nodes method to append additional nodes to subgraphs and Gc-train-to-Gs-train as experimental setup.
5.1.3 Graph Classification
Graph Classification has been conducted on the PROTEINS and AIDS dataset Morris et al. (2020); we randomly split the dataset into a $2:1:1$ ratio of train, validation, and test.
5.1.4 Graph Regression
We perform graph regression on the QM9 Wu et al. (2017), and a subset of ZINC Gómez-Bombarelli et al. (2016) datasets, using the same splits as in graph classification. Gilmer et al. (2017) grouped the $19$ properties of each molecule in QM9 dataset into four broad categories. We predict one property from each of the four broad categories. The four specific properties are the following: the dipole moment ($\mu$), the gap between $\epsilon_{\mathrm{HOMO}}$ and $\epsilon_{\mathrm{LUMO}}$, zero-point vibration energy ($\mathrm{ZPVE}$), and at $298.15\mathrm{K}$, the atomization energy ($U^{\mathrm{ATOM}}$) of each molecule.
5.2 Result and Analysis
FIT-GNN
Dataset Classical $\mathbf{r=0.3}$ $\mathbf{r=0.5}$
ZINC (Subset) $0.688$ $\mathbf{0.578}$ $0.657$
QM9 ($\mu$) $0.869$ $\mathbf{0.841}$ $0.948$
QM9 ($\Delta\epsilon$) $1.012$ $\mathbf{0.875}$ $0.969$
QM9 ($\mathrm{ZPVE}$) $1.091$ $\mathbf{0.818}$ $0.840$
QM9 ($U^{\mathrm{ATOM}}$) $1.078$ $0.754$ $\mathbf{0.630}$
Table 6: Graph regression tasks results on ZINC (Subset) and QM9. The metrics shown are in terms of normalized MAE (lower is better) with Gs-train-to-Gs-infer experimental setup. The Cluster Nodes method is used to append additional nodes to subgraphs.
Gc-train-to
-Gs-infer
Gs-train-to
-Gs-infer
Classical $\mathbf{r=0.3}$ $\mathbf{r=0.5}$ $\mathbf{r=0.3}$ $\mathbf{r=0.5}$
AIDS $0.792$ $0.810$ $0.836$ $0.812$ $\mathbf{0.843}$
PROTEINS $0.623$ $0.636$ $0.681$ $0.812$ $\mathbf{0.821}$
Table 7: Results for the graph classification task on AIDS and PROTEINS are presented. The metrics reflect accuracy; higher values indicate better performance. The table also compares two experimental setups.
Cora Chameleon
Coarsening Method $\mathbf{r=0.1}$ $\mathbf{r=0.3}$ $\mathbf{r=0.1}$ $\mathbf{r=0.3}$
variation_neighborhoods $0.821\pm 0.003$ $0.803\pm 0.005$ $\mathbf{0.465\pm 0.005}$ $\mathbf{0.481\pm 0.006}$
algebraic_JC $\mathbf{0.827\pm 0.006}$ $\mathbf{0.804\pm 0.003}$ $0.544\pm 0.005$ $0.504\pm 0.003$
kron $0.758\pm 0.004$ $0.800\pm 0.002$ $0.571\pm 0.004$ $0.580\pm 0.013$
heavy_edge $0.736\pm 0.012$ $0.773\pm 0.006$ $0.572\pm 0.002$ $0.531\pm 0.002$
variation_edges $0.484\pm 0.006$ $0.471\pm 0.012$ $0.513\pm 0.003$ $0.542\pm 0.005$
variation_cliques $0.751\pm 0.012$ $0.801\pm 0.009$ $0.674\pm 0.009$ $0.679\pm 0.008$
Table 8: Results comparing various coarsening methods on Cora and Chameleon datasets and the metric for each dataset are accuracy (higher the better) and normalized MAE (lower the better) respectively.
Table 3 shows the node regression error and standard deviation on coarsening ratios $0.1$ and $0.3$. All the experiments for node regression are shown with Cluster node method. With a higher coarsening ratio, more subgraphs are created. As the coarsening ratio increases, the size of each subgraph also reduces, and important neighboring node information gets lost. Hence, the performance decreases with an increase in the coarsening ratio.
Table 4 shows the results for node classification accuracy and standard deviation on different coarsening ratios on the Cora dataset with Cluster Nodes approach. With coarsening ratios $0.1$ and $0.3$, the observed performance is better than other coarsening ratios, and the inference time taken is also less. This motivates us to experiment on other datasets with lower coarsening ratios. Table 5 shows the results for node classification on different datasets using coarsening ratios $0.1$ and $0.3$. From the results, it can be observed that the accuracy is comparable to the baseline. Since experimenting on the full OGBN-Products dataset is not feasible due to memory constraints, we create a proxy graph by taking the maximum-sized communities with the number of nodes that sum up to $165000$. For the subset of the OGBN-Products graph, the performance is better with the baseline even with $0.56\%$ training nodes. For larger datasets like Pubmed and Physics Coauthor, a lower coarsening ratio doesn’t perform better than the baseline.
Table 6 shows the results for the graph regression task on ZINC (Subset) and QM9 dataset. All the results show that the FIT-GNN model’s performance is comparable to the baselines. It is also observed that a lower coarsening ratio yields better loss, which implies that the model performs better on molecular graphs when the subgraph size is high. It implies that finer subgraphs lose out global information about the molecule, which is necessary for the prediction.
Table 7 displays graph classification results using two experimental setups across two datasets. Training on coarsened graphs outperforms the baseline but discards important information. Conversely, training on subgraphs significantly improves predictions compared to the baseline.
PROTEINS ZINC (subset)
Coarsening Method $\mathbf{r=0.3}$ $\mathbf{r=0.5}$ $\mathbf{r=0.3}$ $\mathbf{r=0.5}$
variation_neighborhoods $0.652$ $\mathbf{0.739}$ $\mathbf{0.629}$ $0.859$
algebraic_JC $\mathbf{0.783}$ $0.478$ $0.823$ $0.810$
kron $0.522$ $0.652$ $0.663$ $1.201$
heavy_edge $0.565$ $0.609$ $2.114$ $1.227$
variation_edges $0.652$ $0.565$ $1.548$ $1.883$
variation_cliques $0.696$ $0.652$ $0.689$ $\mathbf{0.742}$
Table 9: Results comparing various coarsening methods on PROTEINS and ZINC (subset) datasets. The metric for the PROTEINS dataset is accuracy (higher the better), and for the ZINC (subset) is normalized MAE (lower the better)
Table 8 and Table 9 show the comparison of performance with different coarsening algorithms used by Huang et al. (2021). We find that variation_neighborhoods is more consistent among all other coarsening algorithms, therefore we use it to report other results also.
5.3 Inference Time and Memory Comparison
FIT-GNN
Dataset Classical $\mathbf{r=0.1}$ $\mathbf{r=0.3}$
Chameleon $0.0027$ $0.0016$ $\mathbf{0.0014}$
Squirrel $0.0081$ $0.0017$ $\mathbf{0.0014}$
Crocodile $0.0070$ $0.0015$ $\mathbf{0.0015}$
Cora $0.0026$ $\mathbf{0.0019}$ $0.0020$
Citeseer $0.0031$ $\mathbf{0.0018}$ $0.0019$
Pubmed $0.0042$ $0.0019$ $\mathbf{0.0018}$
DBLP $0.0063$ $0.0020$ $\mathbf{0.0018}$
Physics Coauthor $0.0252$ $0.0020$ $\mathbf{0.0017}$
OGBN-Products $0.1762$ $0.0017$ $\mathbf{0.0016}$
Table 10: Inference time (sec) comparison for a single node using standard GNN as a baseline and FIT-GNN with two different coarsening ratios. Lower is better. Here, we used Cluster Nodes.
Figure 3: Comparison of inference time memory consumption (in MB) for Coauthor Physics Dataset between baseline and FIT-GNN
FIT-GNN
Dastaset Classical $\mathbf{r=0.3}$ $\mathbf{r=0.5}$
ZINC (subset) $\mathbf{0.00184}$ $\mathbf{0.00184}$ $0.00190$
QM9 $\mathbf{0.00173}$ $0.00180$ $0.00191$
AIDS $0.00163$ $\mathbf{0.00155}$ $0.00163$
PROTEINS $0.00165$ $\mathbf{0.00160}$ $0.00163$
Table 11: Inference time (sec) comparison for the graph-level task. We use the Cluster Nodes method to add additional nodes to each subgraph and Gc-train-to-Gc-infer experiment setup
As mentioned in Section 4.3, our inference time is lesser than the standard GNN if certain conditions are met. Therefore, we empirically show reduction of time and space by our FIT-GNN model during inference. For recording the inference time, we use the Python time package to calculate the difference in time before and after the inference step.
In Table 10, we show the average time to predict for $1000$ nodes in the baseline and FIT-GNN models for node regression and classification tasks. The baseline model processes the entire graph, increasing inference time, especially for large graphs. In contrast, the FIT-GNN model only requires the relevant subgraph, resulting in faster predictions. For larger datasets like the OGBN-Products, we see up to $\mathbf{100\times}$ reduction in inference time. Table 17 in the Appendix shows the detailed comparison of memory consumptions for different datasets for both Extra Nodes and Cluster Nodes method along with the baseline memory, which highlights that the FIT-GNN model uses up to $\mathbf{70\times}$ less memory than the baseline.
Table 11 compares the inference time of graph classification and graph regression task. We randomly select $1000$ graphs from the test split and infer them for baseline and FIT-GNN. We also append additional nodes to each subgraph using Cluster Nodes method. The table shows how our method is comparable and sometimes faster than the baseline. Also, we observe that the inference time increases with a higher coarsening ratio. Because, with a higher coarsening ratio, the number of nodes in $G^{\prime}$ increases, resulting in more edges. In table 16 in the Appendix, we show the estimated $\phi_{\textup{max}}$ value from the feature dimension of the nodes and the coarsening ratio. Combinations with negative numbers indicate the FIT-GNN-based approach won’t reduce the inference time.
Overall, the inference time and memory for all the tasks mentioned are drastically less than the classical approach while maintaining the performance mentioned in the previous section.
6 Conclusion
In this paper, we have focused on inference time and memory and presented a new way to utilize existing graph coarsening algorithms for GNNs. We have provided theoretical insight corresponding to the number of nodes in the graph for which the FIT-GNN model reduces the time and space complexity. Empirically, we have shown that our method is comparable to the uncoarsened baseline while being orders of magnitude faster in terms of the inference time and consuming a fraction of the memory. A few possible future directions involve studying directed and weighted graphs, focusing on the theoretical connections between Extra Nodes and Cluster Nodes.
Appendix A Dataset Description
Dataset
Number of
Graphs
Average
Nodes
Average
Edges
Features Classes
PROTEINS $1113$ $19$ $72$ $3$ $2$
AIDS $2000$ $7$ $16$ $38$ $2$
Table 12: Summary of datasets used for Graph classification Dataset
Dataset
Number of
Graphs
Average
Nodes
Average
Edges
Features
Number of
Targets
QM9 $130831$ $8$ $18$ $11$ $19$
ZINC (subset) $10000$ $11$ $25$ $1$ $1$
Table 13: Summary of datasets used for Graph regression Dataset
Dataset Nodes Edges Features Classes
Cora $2708$ $5278$ $1433$ $7$
Citeseer $3327$ $4552$ $3703$ $6$
Pubmed $19717$ $44324$ $500$ $3$
DBLP $17716$ $52867$ $1639$ $4$
Physics
Coauthor
$34493$ $247962$ $8415$ $5$
OGBN-Products $165000$ $4340428$ $100$ $47$
Table 14: Summary of datasets used for Node classification Dataset
Dataset Nodes Edges Features
Number of
Targets
Chameleon $2277$ $31396$ $128$ $1$
Squirrel $5201$ $198423$ $128$ $1$
Crocodile $11631$ $170845$ $128$ $1$
Table 15: Summary of datasets used for Node regression Dataset
Dataset $\mathbf{r=0.1}$ $\mathbf{r=0.3}$ $\mathbf{r=0.5}$ $\mathbf{r=0.7}$
PROTEINS $6.999$ $0.333$ $-1.000$ $-1.571$
AIDS $322.000$ $82.000$ $34.000$ $13.429$
QM9 $79.000$ $19.000$ $7.000$ $1.857$
ZINC (subset) $-11.000$ $-4.333$ $-3.000$ $-2.429$
Table 16: Calculated $\phi_{\textup{max}}$ values for different datasets and coarsening ratios as per Lemma 1.
Appendix B More on Extra Nodes and Cluster Nodes
B.1 Extra Nodes
There are $3$ sets of nodes in $G_{i}$.
  • Nodes with 1-hop and 2-hop neighbours in $G_{i}$. Let us call this set $S_{1}$
  • Nodes with 1-hop neighbours in $G_{i}$ but $\exists v$ in 2-hop neighbourhood that is not in $G_{i}$. Let us call this set $S_{2}$
  • Node where $\exists v$ in 1-hop and 2-hop neighbourhood that is not in $G_{i}$. Let us call this set $S_{3}$.
Let $\mathcal{I}^{2}_{i}$ be the number of nodes whose information doesn’t get passed on after $2$ layer of GNN for $G_{i}$.
$\mathcal{I}^{2}_{i}=\left|\bigcup_{v\in S_{3}}\mathcal{N}_{2}(v)-V(G_{i})\right|$
When we use Extra Nodes, the information loss can be written as follows:
$\mathcal{I}^{2}_{i}=\left|\bigcup_{v\in S_{3}}\mathcal{N}_{2}(v)-\mathcal{E}_{G_{i}}\right|$
The above entity will depend on the density of the subgraphs formed and the number of connections each subgraph shares with each other. An algorithm with an objective to reduce this entity for all subgraphs will lose the least amount of information when Extra Node method is used.
B.2 Cluster Nodes
Given a partition matrix $P$, the features of the coarsened node are $X^{\prime}=P^{T}X$. Given a normalized partition matrix, the features of a node $v^{\prime}_{i}\in G^{\prime}$ is the degree-weighted average of the features of nodes in $C_{i}$. This is one of the functions $f$ used to create the features of the cluster node from $C_{i}$.
Previously, according to Lemma 1, there is no information loss when using the $1$ layer of GNN and Extra Nodes method. It was also easy to quantify in terms of the number of nodes. However, it is different for Cluster Nodes. Let us discuss the issues first.
$\bullet$ Only a weighted version of node information is shared with the subgraph. Suppose $v_{c},v_{d}\in G_{i}$ is connected to $v_{a},v_{b}\in G_{j}$. Then the information contributed by these nodes is $\frac{d(v_{a})x_{a}+d(v_{b})x_{b}}{\sum_{p}d(v_{p})}$. Here $d(v_{p})$ represents the degree of node $v_{p}$, and $x_{p}$ represents the feature of node $v_{p}$.
$\bullet$ Other node information will also be shared which is $\frac{\sum_{p\neq v_{a},v_{b}}d(v_{p})x_{p}}{\sum_{p}d(v_{p})}$. This will capture further dependencies.
The performance of Cluster Node will depend on some distance or similarity metric between $x_{c},x_{d}$ and $f(C_{j})$.
Appendix C Experiment Details (Parameters and Device Configuration)
For node classification and node regression tasks, we use Adam Optimizer with a learning rate of $0.01$ and $L_{2}$ regularization with $0.0005$ weight. We use Adam Optimizer with a learning rate of $0.0001$ and $L_{2}$ regularization with $0.0005$ weight for graph level tasks. For both coarsened graph and set of subgraph based model, we set epochs to 300 and the number of layers of GCN to 2 for training. We set the hidden dimensions for each layer of GCN to $512$.
The device configurations are Intel(R) Xeon(R) Gold 5120 CPU @ $2.20$GHz, $256$GB RAM, NVIDIA A100 $40$GB GPU. We use Pytorch Geometric to to train our models.
Appendix D Results
This section shows the memory consumption for different datasets with different methods. Also, we show how the FIT-GNN model performs with different methods of appending additional nodes to subgraphs at different coarsening ratios.
FIT-GNN
Dataset Split $\mathbf{r=0.1}$ $\mathbf{r=0.3}$ $\mathbf{r=0.5}$ $\mathbf{r=0.7}$ Baseline
Cora cluster $1.310$ $0.648$ $0.694$ $0.868$ $15.743$
extra $2.072$ $1.134$ $1.017$ $1.017$
Citeseer cluster $31.559$ $2.148$ $1.253$ $1.686$ $49.488$
extra $31.559$ $3.114$ $1.788$ $1.788$
Pubmed cluster $0.613$ $0.569$ $0.571$ $0.573$ $41.227$
extra $0.798$ $0.590$ $0.590$ $0.590$
DBLP cluster $6.067$ $2.940$ $1.847$ $1.966$ $118.121$
extra $13.907$ $4.663$ $2.649$ $2.274$
Physics Coauthor cluster $24.025$ $13.088$ $13.388$ $14.353$ $1169.521$
extra $63.146$ $22.434$ $22.434$ $22.434$
Chameleon cluster $0.212$ $0.246$ $0.291$ $0.438$ $2.198$
extra $0.653$ $0.653$ $0.653$ $0.592$
Crocodile cluster $1.183$ $1.188$ $1.735$ $1.866$ $11.562$
extra $4.368$ $4.368$ $4.368$ $3.134$
Squirrel cluster $1.507$ $1.607$ $1.903$ $3.009$ $9.075$
extra $7.275$ $7.275$ $7.275$ $6.953$
OGBN-Products cluster $6.047$ $5.613$ $8.636$ $208.029$
Table 17: Summary of memory consumption of datasets used for node-level tasks. All units are in MegaBytes (MB)
r Extra Nodes Cluster Nodes
$0.1$ $0.798\pm 0.008$ $\mathbf{0.824\pm 0.003}$
$0.3$ $0.798\pm 0.008$ $\mathbf{0.807\pm 0.004}$
$0.5$ $0.807\pm 0.004$ $\mathbf{0.814\pm 0.004}$
$0.7$ $0.797\pm 0.007$ $\mathbf{0.808\pm 0.006}$
Table 18: Average of top 10 accuracies (higher the better) of node classification task on Cora dataset with Gs-train-to-Gs-infer experimental setup. The table also demonstrates the superior performance of the Cluster Nodes method over Extra Nodes.
r Extra Nodes Cluster Nodes
0.1 $0.504\pm 0.004$ $\mathbf{0.460\pm 0.003}$
0.3 $0.579\pm 0.006$ $\mathbf{0.499\pm 0.006}$
0.5 $0.585\pm 0.004$ $\mathbf{0.505\pm 0.004}$
0.7 $\mathbf{0.504\pm 0.006}$ $0.560\pm 0.004$
Table 19: Average of best 10 losses (lesser the better) of node regression task on Chameleon dataset with Gs-train-to-Gs-infer experimental setup. The table also demonstrates the better performance of the Cluster Nodes method over Extra Nodes.
Dataset Coarsening Ratio Top 10 Loss
Chameleon Baseline $0.844\pm 0.000$
$0.1$ $\mathbf{0.460\pm 0.002}$
$0.3$ $0.499\pm 0.006$
$0.5$ $0.587\pm 0.000$
$0.7$ $0.591\pm 0.080$
Crocodile Baseline $0.853\pm 0.000$
$0.1$ $\mathbf{0.362\pm 0.002}$
$0.3$ $0.364\pm 0.001$
$0.5$ $0.375\pm 0.001$
$0.7$ $0.384\pm 0.002$
Squirrel Baseline $0.809\pm 0.000$
$0.1$ $\mathbf{0.648\pm 0.002}$
$0.3$ $0.672\pm 0.003$
$0.5$ $0.729\pm 0.004$
$0.7$ $0.694\pm 0.004$
Table 20: Node Regression Results with Gs-train-to-Gs-infer experimental setup.