### v4

# FIT-GNN: Faster Inference Time for GNNs that ‘FIT’ in Memory Using Coarsening

Abstract
Scalability of Graph Neural Networks (GNNs) remains a significant challenge. To tackle this, methods like coarsening, condensation, and computation trees are used to train on a smaller graph, resulting in faster computation. Nonetheless, prior research has not adequately addressed the computational costs during the inference phase. This paper presents a novel approach to improve the scalability of GNNs by reducing computational burden during the inference phase using graph coarsening. We demonstrate two different methods – Extra Nodes and Cluster Nodes. Our study extends the application of graph coarsening for graph-level tasks, including graph classification and graph regression. We conduct extensive experiments on multiple benchmark datasets to evaluate the performance of our approach. Our results show that the proposed method achieves orders of magnitude improvements in single-node inference time compared to traditional approaches. Furthermore, it significantly reduces memory consumption for node and graph classification and regression tasks, enabling efficient training and inference on low-resource devices where conventional methods are impractical. Notably, these computational advantages are achieved while maintaining competitive performance relative to baseline models.
1 Introduction
Graph Neural Networks (GNNs) have demonstrated remarkable versatility and modeling capabilities. However, several challenges still hinder their widespread applicability, with scalability being the most significant concern. This scalability issue affects both the training and inference phases. Previous research has sought to mitigate the high training computation cost by using three broad categories of methods: 1. Training every iteration on a sampled subgraph 2. Training on a coarsened graph generated from a coarsening algorithm 3. Training on a synthetic graph that mimics the original graph . Unfortunately, these smaller graphs can only be utilized for training purposes and have not alleviated the computational burden during inference since the method mentioned above requires the whole graph during the inference phase. Consequently, for larger graphs, not only do inference times increase, but memory consumption may also exceed limitations.
To mitigate the high computational cost associated with the inference phase, we decompose the original graph $G$ into a collection of subgraphs $\mathcal{G}_{s}=\{G_{1},G_{2},\dots,G_{k}\}$. This transformation yields substantial improvements in inference efficiency, achieving up to a $\mathbf{100\times}$ speedup for single-node prediction. Moreover, it significantly reduces memory consumption by as much as $\mathbf{100\times}$ —thereby enabling scalable inference on large-scale graphs comprising millions of edges. We provide a theoretical characterization of the upper bound on the coarsening ratio under which our method’s time and space complexity remain superior to that of the standard approach. Notably, we observe that conventional methods often fail to perform inference on large graphs due to memory constraints, whereas our approach remains effective and scalable in such scenarios.
While partitioning the graphs into subgraphs seems to be a naive transformation, it comes with the cost of information loss corresponding to the edges removed. We overcome this loss by appending additional nodes in the subgraphs using two different methods – Extra Nodes and Cluster Nodes that respectively append neighboring nodes or representational nodes corresponding to neighboring clusters. These additional nodes resulted in equivalent or better performance in terms of accuracy and error compared to baselines. Our contributions are summarized as follows:
  1. 1.
    We reduce inference time and memory by applying graph coarsening to partition graphs into subgraphs, mitigating information loss through Extra Node and Cluster Node strategies.
  2. 2.
    We provide a theoretical analysis of our approach, establishing a lower inference computational complexity compared to baselines.
  3. 3.
    Our method scales well to extremely large graphs that baselines fail to process, while consistently delivering strong performance on 13 real-world datasets.
Figure 1: The Figure shows the overall pipeline of our proposed method and its comparison with traditional training and inference of GNNs. This pipeline is made for node-level tasks.
2 Related Work
The aim of reducing computational costs has been approached through various methods. Research by 10.1145/3292500.3330925 and Zeng2020GraphSAINT: uses a subgraph sampling technique, where a small subgraph is sampled at each training iteration for GNN training. Other methods, including layer-wise sampling and mini-batch training, have been explored by chen2018stochastic; chen2018fastgcn; 10.1145/3394486.3403192; DBLP:journals/corr/abs-1911-07323. Additionally, DBLP:journals/corr/abs-2007-02817 introduced a coarsening algorithm based on Schur complements to create node embeddings in large graphs, while 10.1145/3447548.3467256 [SGGC] applied GNNs to coarsened graphs, reducing nodes and edges to lower training time. pmlr-v202-kumar23a; joly2024graph; NEURIPS2024_733209a1; 10.1145/3589335.3651920 also proposed different coarsening algorithms for scalable training approaches. jin2021graph [GCOND], wang2024fast [GC-SNTK] presented algorithms to generate a synthetic graph that mimics the original graph. However, their approach generates model-specific synthetic graphs that require training on the whole dataset, which defeats the purpose of reducing computational burden. gupta2025bonsai [BONSAI] addressed this issue by looking into computational trees and training on just the relevant computational trees. However, these approaches do not decrease inference time since the inference phase takes the whole graph as input. DBLP:journals/corr/abs-2202-00075 shifted the training process to subgraph-level, using coarsening algorithms to split the graph into $k$-subgraphs and recommending a multi-device training approach. In the graph-level task, Jin_2022 [DOSCOND] and 10.1145/3580305.3599398 [KIDD] tackled the graph classification problem following up on the approach of GCOND and GC-SNTK. However, the graph regression problem was not explored in this direction.
3 Preliminaries
3.1 Graph Notations
Given an undirected graph $G=(V,E,X,W)$, $V$ is the vertex set, $E$ is the edge set, $X\in\mathbb{R}^{n\times d}$ is the feature matrix, and $W$ is the edge weight matrix. Let $|V|=n$ be the number of nodes and $|E|=m$ be the number of edges. Let $A\in\{0,1\}^{n\times n}$ be the adjacency matrix of $G$, $d_{i}$ to be the degree of the node $v_{i}$. $D\in\mathbb{R}^{n\times n}$ is a diagonal degree matrix with $i$th diagonal element as $d_{i}$. We denote $\mathcal{N}_{j}(v_{i})$ as the $j$-hop neighbourhood of node $v_{i}\in V$.
3.2 Graph Neural Network
Graph Neural Network is a neural network designed to work for graph-structured data. The representation of each node in the graph is updated recursively by aggregation and transforming the node representations of its neighbors. kipf2017semisupervised introduced Graph Convolutional Network (GCN) as follows:
$X^{(l+1)}=\sigma(\tilde{D}^{-\frac{1}{2}}\tilde{A}\tilde{D}^{-\frac{1}{2}}X^{(l)}\mathcal{W}^{(l)})$ (1)
where $X^{(l)}$ is the node representation after $l$ layers, $\tilde{A}=A+I$, $\tilde{D}=D+I$, $I$ is the identity matrix of same dimension as $A$. $\mathcal{W}^{(l)}$ is the learnable parameter and $\sigma$ is a non-linear activation function.
3.3 Graph Coarsening and Graph Partitioning
loukas2018graphreductionspectralcut discussed multiple coarsening algorithms which create a coarsened graph $G^{\prime}=(V^{\prime},E^{\prime},X^{\prime},W^{\prime})$ from a given graph $G=(V,E,X,W)$. $X^{\prime}$ is the coarsened node representation. We refer the vertex set $V^{\prime}=\{v^{\prime}_{1},v^{\prime}_{2},\dots,v^{\prime}_{k}\}$ of $G^{\prime}$ as coarsened nodes. Given a coarsening ratio $r\in[0,1]$, we have $n\times r, kk\in\mathbb{Z}$. We can interpret it as creating $k$ disjoint partitions, $C_{1},C_{2},\dots,C_{k}$, from a graph of $n$ nodes. Mathematically, we create a partition matrix $P\in\{0,1\}^{n\times k}$, where $P_{ij}=1$ if and only if node $v_{i}\in C_{j}$. $A^{\prime}=P^{T}AP$ is the adjacency matrix of $G^{\prime}$. Similarly, the corresponding degree matrix $D^{\prime}=P^{T}DP$. SGGC (10.1145/3447548.3467256) used normalized partition matrix $\mathcal{P}=PC^{-\frac{1}{2}}$, where $C$ is defined as a diagonal matrix with diagonal entries $C_{jj}=|C_{j}|,j=1,2,\dots,k$.
Along with $G^{\prime}$, we create a set of disjoint subgraphs $\mathcal{G}_{s}=\{G_{1},G_{2},$ $\dots,G_{k}\}$ corresponding to the partitions $C_{1},C_{2},\dots,C_{k}$. The number of nodes in each partition $C_{i}$ is denoted by $n_{i}$. Each subgraph $G_{i}$ is the induced subgraph of $G$ formed by the nodes in $C_{i}$. $A_{i}$ and $D_{i}$ are the adjacency matrix and degree matrix of $G_{i}$ respectively.
[…]
images/img2
Figure 2: Figure showing the comparison between the Extra Node Method and the Cluster Node Method of appending additional nodes in $G_{1},G_{2},G_{3}$.
4 Methodology
In SGGC, a graph $G$ is reduced to $G^{\prime}$ using a partition matrix $P$ generated from the coarsening algorithm. Here $X^{\prime}=P^{T}X$. The labels for the coarsened graph are $Y^{\prime}=\arg\max(P^{T}Y)$, where $Y$ is the label matrix of $G$ storing the one-hot encoding of the label of each node for the node classification task. Training is carried out on the coarsened graph $G^{\prime}$ using $Y^{\prime}$ as the label information. While this approach is novel, essential label information is lost. For instance, suppose a node $v^{\prime}\in G^{\prime}$ is a combination of $v_{i},v_{j},v_{k}\in G$ where $v_{i},v_{j},v_{k}$ are test nodes. Let us assume that the classes of $v_{i},v_{j},v_{k}$ are $0,0,2$ respectively. $\arg\max$ will take the majority label and assign class 0 to $v^{\prime}$. Hence, the model will be trained to predict only the class 0, leading to the discarding of the quantification of the model’s performance on predicting less represented nodes.
Similarly, to explore another node-level task, such as node regression, aggregation functions such as the mean of regression targets wouldn’t be appropriate because they lose the variance of the targets. To mitigate label information loss, we use original label information without any aggregation function like $\arg\max$ or mean, thus motivating us to follow a subgraph-level training process.
However, partitioning the graph into subgraphs leads to a loss of neighborhood information at the periphery of subgraphs. To tackle this, we append nodes in two ways after partitioning:
  • Extra Nodes: Adding the $1$-hop neighbouring nodes, namely Extra Nodes, in each subgraph $G_{i}$ denoted by $\mathcal{E}_{G_{i}}$. DBLP:journals/corr/abs-2202-00075 introduced this approach to reduce information loss after partitioning. We add a unit weight edge if two nodes in $\mathcal{E}_{G_{i}}$ are connected in $G$. Node embedding of $v\in\mathcal{E}_{G_{i}}$ is $x_{v}\in X$.
    $\mathcal{E}_{G_{i}}=\bigcup_{v\in G_{i}}\left\{u\in V:u\in\mathcal{N}_{1}(v)\wedge u\notin G_{i}\right\}$ (2)
  • Cluster Nodes: Although adding $1$-hop neighboring nodes in the subgraph helps recover lost information after partition, for multi-layered GNN models, information loss will persist. Instead of adding the neighboring nodes, we will add Cluster Nodes, which represent the neighboring clusters, denoted by $\mathcal{C}_{G_{i}}$. liu2024scalable introduced this method to add nodes in the graph instead of adding to separate subgraphs. They also added cross-cluster edges among these Cluster Nodes and proposed a subgraph sampling-based method on the modified graph. In our work, we add cross-cluster edges. For each node $t\in\mathcal{C}_{G_{i}}$, the node embedding $x_{t}=X^{\prime}_{t}$.
    $\mathcal{C}_{G_{i}}=\bigcup_{v\in\mathcal{E}_{G_{i}}}\left\{t:P_{v,t}\neq 0\right\}$ (3)
Figure 2 illustrates appending these additional nodes using these approaches. The dashed line shows which nodes are added as part of these additional nodes. For subgraph-level training, specifically for node-level tasks, the newly appended nodes do not contribute to the weight update as the loss is not backpropagated based on the predictions of these nodes. This is done since the appended nodes might contain information from the test nodes. Therefore, a boolean array $mask_{i}$ is created for each subgraph $G_{i}$ such that $True$ is set for nodes that actually belong to the subgraph $G_{i}$ (not as an Extra Node or Cluster Node) and are training nodes; otherwise, $False$.
Lemma 4.1.
Models with $1$ layer of GNN cannot distinguish between $G$ and $\mathcal{G}_{s}$ when Extra Nodes method is used. (Proof in Appendix A.1)
Using the Cluster Nodes approach to append additional nodes to subgraphs is better for two reasons:
  • From Equation 2 and 3, we can say that $\sum_{i}|\mathcal{E}_{G_{i}}|\geq\sum_{i}|\mathcal{C}_{G_{i}}|$. This is because, using Extra Nodes, we add all the neighboring nodes. However, using Cluster Nodes, we only add one representative node for each neighboring cluster. Hence, the time taken to train and infer these modified subgraphs with Cluster Nodes is less than or equal to that with Extra Nodes.
  • Lemma 4.1 shows that the Extra Nodes approach reduces the information loss due to partitioning for $1$ Layer of GNN. However, with multiple layers of GNN, longer dependencies are not captured. In contrast, the Cluster Nodes approach overcomes this by computing a single Cluster Node’s representation as a combination of features of all the nodes present in the corresponding cluster. This results in sharing further node information in $1$-hop. Additionally, the transfer of information from one cluster to another is captured.
Algorithm 1 Training GNN (Train on $\mathcal{G}_{s}$)
1:$G=(V,E,X)$; Labels $Y$; Node-Model $M$; Loss $\ell$; Number of Layers $L$
2:Apply a coarsening algorithm on $G$, and output a normalized partition matrix $P$.
3:Construct $\mathcal{G}_{s}=\{G_{1},G_{2},\dots,G_{k}\}$ using $P$;
4:if Append additional nodes then
[…]
6:end if
7:Construct mask for each subgraph $Mask=\{mask_{1},mask_{2},\dots,mask_{k}\}$;
8:Construct feature matrix $\{X_{1},X_{2},\dots,X_{k}\}$ for $\mathcal{G}_{s}$;
9:Construct label matrix $\{Y_{1},Y_{2},\dots,Y_{k}\}$ for $\mathcal{G}_{s}$ ;
[…]
11:for each subgraph $G_{i}$ in $\mathcal{G}_{s}$ do
12:  $O_{i}=M(L,A_{i},D_{i},X_{i})$; $\quad\#$ Algorithm 4
13:  $O.append(O_{i}[mask_{i}])$;
14:  $Y.append(Y_{i}[mask_{i}])$;
15:end for
[…]
17:Train $M$ to minimize LossLoss;
4.1 Node-level Task
We use Algorithm 4 with $L$ number of layers to create a standard GNN model $M$ for node-level tasks. While we use Algorithm 3 to train a model on the coarsened graph $G^{\prime}$, we propose Algorithm 1 to train a model on the set of subgraphs $\mathcal{G}_{s}$. Algorithm 3 utilizes the partition matrix $P$ generated by the coarsening algorithm to construct and train on the coarsened graph $G^{\prime}$ and the coarsened label information $Y^{\prime}$. On the other hand, our proposed Algorithm 1 uses the same partition matrix $P$ to construct and train on subgraphs $\mathcal{G}_{s}$ using original label information $Y$. For the node classification task, $\arg\max$ is used for creating labels for $G^{\prime}$ and CrossEntropy as a loss function. For the node regression task, we do not create $G^{\prime}$. Mean Absolute Error (MAE) is used as the loss function.
4.2 Graph-level Task
For a given graph $G\in\mathcal{D}$, we create $G^{\prime}$ and $\mathcal{G}_{s}=\{G_{1},G_{2},$ $\dots,G_{k}\}$. Since $G^{\prime}$ is a single graph and $\mathcal{G}_{s}$ is a set of subgraphs, we use Algorithm 5 to create a model to train on $G^{\prime}$ and Algorithm 2 to create a model to train on $\mathcal{G}_{s}$. We use the CrossEntropy loss function for the graph classification task and the Mean Absolute Error(MAE) for the graph regression task.
Algorithm 2 Graph-Model-$\mathcal{G}_{s}(\mathcal{G}_{s},L)$
[…]
3:for each subgraph $G_{i}\in\mathcal{G}_{s}$ do
4:  for $j=1$ to $L$ do
[…]
6:  end for
[…]
8:end for
[…]
11:return $Z$
4.3 Time and Space Complexity
Table 1: Training and inference time and space complexity of FIT-GNN compared to classical and SGGC approaches.
(a) Time Complexity
Train Inference (Full Graph)
Classical $nd^{2}+n^{2}d$ $nd^{2}+n^{2}d$
SGGC $kd^{2}+k^{2}d$ $nd^{2}+n^{2}d$
FIT-GNN $\begin{matrix}kd^{2}+k^{2}d\\ +\sum_{i=1}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]\end{matrix}$ $\sum_{i=1}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]$
(b) Space Complexity
Train Inference
Classical $n^{2}+nd+d^{2}$ $n^{2}+nd+d^{2}$
SGGC $k^{2}+kd+d^{2}$ $n^{2}+nd+d^{2}$
FIT-GNN $\begin{matrix}k^{2}+kd+d^{2}\\ +\max_{i=1}[\bar{n}_{i}^{2}+\bar{n}_{i}d]\end{matrix}$ $d^{2}+\max_{i=1}[\bar{n}_{i}^{2}+\bar{n}_{i}d]$
In Equation 1, the matrix dimensions are as follows: $\tilde{A}$ has dimensions $(n\times n)$, $X^{(l)}$ is of dimensions $(n\times d)$, and $\mathcal{W}^{(l)}$ has dimensions $(d\times d)$. Time complexity for one layer of GNN computation is $\mathcal{O}(n^{2}d+nd^{2})$. If we take $L$ layers, then the total time is $\mathcal{O}(Ln^{2}d+Lnd^{2})$. The space complexity is $\mathcal{O}(n^{2}+Lnd+Ld^{2})$. When we compute on a sparse graph, the time complexity is $\mathcal{O}(m+Lnd^{2})$ and space complexity is $\mathcal{O}(m+Lnd+Ld^{2})$. The complexity is based on GCN. For other architectures like GAT veličković2018graphattentionnetworks, the complexity changes.
SGGC improved the time and space complexity (Table 1) for training the network by reducing the number of nodes from $n$ to $k$. However, the inference time and space complexity remain the same.
Let us compare three different models. One is the classical model, where no coarsening or partitioning is done; the second is SGGC, where training is performed on a smaller graph $G^{\prime}$ and inference on $G$; the third is our approach, FIT-GNN, where both training and inference are performed on $\mathcal{G}_{s}$.
The inference time complexity of our model is $\sum_{i=1}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]$, where $\bar{n}_{i}=n_{i}+\phi_{i}$ and $\phi_{i}$ is the number of additional nodes appended in each subgraph. Table 1 compares different approaches.
Lemma 4.2.
The inference time complexity is at most $\sum_{i=1}^{k}[(n_{i}+\phi_{i})^{2}d+(n_{i}+\phi_{i})d^{2}]\leq n^{2}d+nd^{2}$ if $\mathbb{E}[n_{i}+\phi_{i}]\leq\sqrt{\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})}-\frac{d}{2}$.
Proof.
$\sum_{i=1}^{k}[(n_{i}+\phi_{i})^{2}d+(n_{i}+\phi_{i})d^{2}]=\sum_{i=1}^{k}[n_{i}^{2}d+\phi_{i}^{2}d+2n_{i}\phi_{i}d+n_{i}d^{2}+\phi_{i}d^{2}]$
$=nd^{2}+d\left[\sum^{k}_{i=1}n_{i}^{2}+\sum^{k}_{i=1}\phi_{i}^{2}+2\sum_{i=1}^{k}n_{i}\phi_{i}+d\sum_{i=1}^{k}\phi_{i}\right]$
Now, define $\mathbb{E}[n_{i}]:=\frac{1}{k}\sum_{i=1}^{k}n_{i}=\frac{n}{k}=\frac{1}{r}$ (since $k=nr$) and $\mathbb{E}[\phi_{i}]:=\frac{1}{k}\sum_{i=1}^{k}\phi_{i}$. We also define the variances and covariance,
$\text{Var}(n_{i}):=\frac{1}{k}\sum_{i=1}^{k}n_{i}^{2}-\mathbb{E}[n_{i}]^{2}$ $\qquad\text{Var}(\phi_{i}):=\frac{1}{k}\sum_{i=1}^{k}\phi_{i}^{2}-\mathbb{E}[\phi_{i}]^{2},$
$\text{Cov}(n_{i},\phi_{i})=\text{Cov}$ $:=\frac{1}{k}\sum_{i=1}^{k}\phi_{i}n_{i}-\mathbb{E}[\phi_{i}]\mathbb{E}[n_{i}]$
We can also write,
$\text{Var}(n_{i}+\phi_{i})=\frac{1}{k}\sum_{i=1}^{k}(n_{i}+\phi_{i})^{2}-(\mathbb{E}[n_{i}]+\mathbb{E}[\phi_{i}])^{2}=\text{Var}(n_{i})+\text{Var}(\phi_{i})+2\text{Cov}(n_{i},\phi_{i})$
Using the definitions above, we can say:
$\sum_{i=1}^{k}[(n_{i}+\phi_{i})^{2}d+(n_{i}+\phi_{i})d^{2}]=nd^{2}+d\left[\sum^{k}_{i=1}n_{i}^{2}+\sum^{k}_{i=1}\phi_{i}^{2}+2\sum_{i=1}^{k}n_{i}\phi_{i}+d\sum_{i=1}^{k}\phi_{i}\right]$
$=nd^{2}+d\left[k(\textup{Var}(n_{i})+\mathbb{E}[n_{i}]^{2})+k(\textup{Var}(\phi_{i})+\mathbb{E}[\phi_{i}]^{2})+2k(\textup{Cov}(n_{i},\phi_{i})+\mathbb{E}[n_{i}]\mathbb{E}[\phi_{i}])+dk\mathbb{E}[\phi_{i}]\right]$
$=nd^{2}+nd\left[r\textup{Var}(n_{i})+\frac{1}{r}+r\textup{Var}(\phi_{i})+r\mathbb{E}[\phi_{i}]^{2}+2r\textup{Cov}(n_{i},\phi_{i})+2\mathbb{E}[\phi_{i}]+rd\mathbb{E}[\phi_{i}]\right]$
$=nd^{2}+nd\left[r\mathbb{E}[\phi_{i}]^{2}+(2+rd)\mathbb{E}[\phi_{i}]+r\textup{Var}(n_{i}+\phi_{i})+\frac{1}{r}\right]$
Given, $\mathbb{E}[n_{i}+\phi_{i}]\leq\sqrt{\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})}-\frac{d}{2}\Rightarrow\mathbb{E}[\phi_{i}]\leq\sqrt{\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})}-(\frac{d}{2}+\frac{1}{r})$. Using this, we can say the following:
$\sum_{i=1}^{k}[(n_{i}+$ $\phi_{i})^{2}d+(n_{i}+\phi_{i})d^{2}]=nd^{2}+nd\left[r\mathbb{E}[\phi_{i}]^{2}+(2+rd)\mathbb{E}[\phi_{i}]+r\textup{Var}(n_{i}+\phi_{i})+\frac{1}{r}\right]$
$\leq nd^{2}+nd\left[r\left(\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})+\frac{d^{2}}{4}+\frac{1}{r^{2}}+\frac{d}{r}-2\left(\frac{d}{2}+\frac{1}{r}\right)\sqrt{\Delta}\right)\right$.
$\qquad\qquad\left.+2\sqrt{\Delta}-2\left(\frac{d}{2}+\frac{1}{r}\right)+rd\sqrt{\Delta}-rd\left(\frac{d}{2}+\frac{1}{r}\right)+r\textup{Var}(n_{i}+\phi_{i})+\frac{1}{r}\right]$
$\qquad\qquad\qquad(\textup{where }\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})=\Delta)$
$=nd^{2}+n^{2}d$
Hence, $\sum_{i=1}^{k}[(n_{i}+\phi_{i})^{2}d+(n_{i}+\phi_{i})d^{2}]\leq n^{2}d+nd^{2}$. ∎
Corollary 4.3.
$\mathbb{E}[\phi_{i}]$ has a positive upper bound when $\textup{Var}(n_{i}+\phi_{i})\leq\frac{n}{r}-\frac{1}{r^{2}}$.
Proof.
From Lemma 4.2, we know the upper bound of $\mathbb{E}[\phi_{i}]$ is $\sqrt{\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})}-\left(\frac{d}{2}+\frac{1}{r}\right)$.
If $\sqrt{\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})}-\left(\frac{d}{2}+\frac{1}{r}\right)\geq 0$, then we can say the following:
$\sqrt{\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})}-\left(\frac{d}{2}+\frac{1}{r}\right)\geq 0$
$\Rightarrow\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})\geq\left(\frac{d}{2}+\frac{1}{r}\right)^{2}$
$\Rightarrow\textup{Var}(n_{i}+\phi_{i})\leq\frac{n}{r}-\frac{1}{r^{2}}$
For a given graph, when the above conditions are satisfied, our approach has better inference time and space complexity than other approaches. From Corollary 4.3, we can understand that it is ideal to have similarly sized subgraphs. For more details on time and space complexity, refer to Section C in the Appendix.
5 Experiments
Let training on the coarsened graph $G^{\prime}$ be referred to as Gc-train, subgraph-level training as Gs-train, and subgraph-level inference process as Gs-infer. Once we have constructed $G^{\prime}$ and $\mathcal{G}_{s}$, we train FIT-GNN in 4 different setups:
  • Gc-train-to-Gs-train: Train the GNN model on $G^{\prime}$ as per Algorithm 3, then use the learned weight as an initialization for subgraph-level training and final inference on $\mathcal{G}_{s}$. Our aim was to examine whether pretraining on $G^{\prime}$ (which requires less time) followed by fine-tuning on $\mathcal{G}_{s}$ for fewer epochs can improve performance.
  • Gc-train-to-Gs-infer: Train the GNN model on $G^{\prime}$ as per Algorithm 3, then infer on $\mathcal{G}_{s}$ using the learned weights. We explore whether training solely on the $G^{\prime}$ is sufficient for inference on $\mathcal{G}_{s}$. This setting presents a computationally efficient option, allowing users with limited resources to obtain results close to the best reported.
  • Gs-train-to-Gs-infer: Perform subgraph-level training and inference on $\mathcal{G}_{s}$.
  • Gc-train-to-Gc-infer: Unlike node-level tasks, for graph-level tasks, inference can be done directly on $G^{\prime}$ because the label corresponds to the entire graph and not to individual nodes. Therefore, even with a reduced graph representation, meaningful inference is possible. Therefore, in this setup for graph-level tasks only, we train and infer on $G^{\prime}$.
All four setups can be applied to graph-level tasks. For the node classification task, we apply all setups except Gc-train-to-Gc-infer. For the node regression task, we only perform Gs-train-to-Gs-infer because a coarsened graph is not created for node regression datasets.
Dataset descriptions and details related to hyperparameters, device configuration, and key packages used for the experiments are in Section D and E. Source code: FIT-GNN.
Table 2: Datasets and corresponding splits used in our experiments. ‘c’ denotes the number of classes in node classification datasets. Public split “fixed” from 10.5555/3045390.3045396 is used for Cora, Citeseer, and PubMed. For the “random” split, we use 20 nodes per class for training, 30 per class for validation, and the rest for testing. For QM9, we follow DBLP:journals/corr/GilmerSRVD17 and predict one property from each broad category: dipole moment ($\mu$), HOMO-LUMO gap ($\epsilon_{\mathrm{HOMO}}-\epsilon_{\mathrm{LUMO}}$), zero-point vibration energy (ZPVE), and atomization energy ($U^{\mathrm{ATOM}}$) at $298.15\mathrm{K}$.
Task Dataset (Reference) Train Val Test
Node Regression Chameleon rozemberczki2019multiscale 30% 20% 50%
Crocodile rozemberczki2019multiscale 30% 20% 50%
Squirrel rozemberczki2019multiscale 30% 20% 50%
Node Classification Cora McCallum2000 $20\times c$ $30\times c$ Rest
Citeseer 10.1145/276675.276685 $20\times c$ $30\times c$ Rest
Pubmed Sen_Namata_Bilgic_Getoor_Galligher_Eliassi-Rad_2008 $20\times c$ $30\times c$ Rest
Coauthor Physics DBLP:journals/corr/abs-1811-05868 $20\times c$ $30\times c$ Rest
DBLP 10.1145/1401890.1402008 $20\times c$ $30\times c$ Rest
OGBN-Products 10.5555/3495724.3497579 $20\times c$ $30\times c$ Rest
Graph Regression QM9 DBLP:journals/corr/WuRFGGPLP17 50% 25% 25%
ZINC (Subset) DBLP:journals/corr/Gomez-Bombarelli16 50% 25% 25%
Graph Classification PROTEINS Morris+2020 50% 25% 25%
AIDS Morris+2020 50% 25% 25%
6 Result and Analysis
6.1 Model Performance
Table 3: This table shows the normalized MAE loss (lower is better) for node regression task on different datasets with Gs-train-to-Gs-infer experiment setup and Cluster Nodes method. variation_neighborhoods coarsening algorithm is used.
Methods Models Reduction Ratio (r) Dataset
Chameleon Crocodile Squirel
Full GCN 1.0 $0.852\pm 0.001$ $0.854\pm 0.000$ $0.802\pm 0.000$
GAT 1.0 $0.846\pm 0.002$ $0.850\pm 0.000$ $0.805\pm 0.001$
SAGE 1.0 $0.848\pm 0.002$ $0.850\pm 0.000$ $0.804\pm 0.000$
GIN 1.0 $0.843\pm 0.000$ $0.852\pm 0.000$ $0.796\pm 0.000$
FIT-GNN GCN 0.1 $0.496\pm 0.005$ $0.364\pm 0.001$ $0.663\pm 0.003$
0.3 $0.531\pm 0.006$ $0.369\pm 0.002$ $0.682\pm 0.002$
0.5 $0.531\pm 0.004$ $0.371\pm 0.001$ $0.722\pm 0.002$
0.7 $0.536\pm 0.009$ $0.368\pm 0.003$ $0.696\pm 0.002$
GAT 0.1 $0.582\pm 0.029$ $0.395\pm 0.022$ $0.734\pm 0.019$
0.3 $0.537\pm 0.021$ $0.383\pm 0.020$ $0.724\pm 0.020$
0.5 $0.556\pm 0.018$ $0.386\pm 0.018$ $0.728\pm 0.023$
0.7 $0.576\pm 0.018$ $0.395\pm 0.012$ $0.746\pm 0.020$
SAGE 0.1 $0.506\pm 0.006$ $\mathbf{0.309\pm 0.014}$ $\mathbf{0.589\pm 0.003}$
0.3 $0.489\pm 0.003$ $0.315\pm 0.001$ $0.602\pm 0.003$
0.5 $\mathbf{0.484\pm 0.004}$ $0.317\pm 0.001$ $0.618\pm 0.002$
0.7 $0.539\pm 0.009$ $0.323\pm 0.000$ $0.621\pm 0.004$
GIN 0.1 $0.769\pm 0.101$ $0.722\pm 0.098$ $0.795\pm 0.018$
0.3 $0.808\pm 0.056$ $0.697\pm 0.122$ $0.796\pm 0.004$
0.5 $0.810\pm 0.062$ $0.730\pm 0.091$ $0.786\pm 0.032$
0.7 $0.792\pm 0.089$ $0.779\pm 0.102$ $0.791\pm 0.009$
Table 4: Results for node classification tasks with accuracy as the metric (higher is better). We use the Cluster Nodes method to append additional nodes to subgraphs and Gs-train-to-Gs-infer as experimental setup. variation_neighborhoods coarsening algorithm is used.
Methods Models
Reduction
Ratio (r)
Dataset
Cora Citeseer Pubmed DBLP Physics
Full GCN 1.0 $0.821\pm 0.002$ $0.706\pm 0.002$ $0.768\pm 0.020$ $0.724\pm 0.002$ $\mathbf{0.933\pm 0.000}$
GAT 1.0 $0.809\pm 0.004$ $0.700\pm 0.003$ $0.740\pm 0.004$ $0.713\pm 0.004$ $0.909\pm 0.008$
SGGC GCN 0.3 $0.808\pm 0.003$ $0.700\pm 0.001$ $0.773\pm 0.002$ $0.737\pm 0.011$ $0.928\pm 0.003$
0.5 $0.808\pm 0.001$ $0.716\pm 0.002$ $\mathbf{0.793\pm 0.002}$ $0.733\pm 0.008$ $0.931\pm 0.004$
GAT 0.3 $0.686\pm 0.010$ $0.635\pm 0.021$ $0.773\pm 0.002$ $0.626\pm 0.028$ $0.870\pm 0.016$
0.5 $0.641\pm 0.026$ $0.635\pm 0.024$ $0.683\pm 0.024$ $0.604\pm 0.047$ $0.842\pm 0.018$
GCOND GCN 0.3 $0.806\pm 0.003$ $0.722\pm 0.001$ $0.779\pm 0.002$ $0.748\pm 0.001$ $0.876\pm 0.011$
0.5 $0.808\pm 0.004$ $\mathbf{0.730\pm 0.007}$ $0.773\pm 0.004$ $0.755\pm 0.002$ $0.804\pm 0.005$
GAT 0.3 $0.549\pm 0.080$ $0.628\pm 0.095$ $0.370\pm 0.079$ $0.481\pm 0.034$ OOM
0.5 $0.712\pm 0.059$ $0.501\pm 0.085$ $0.409\pm 0.044$ $0.464\pm 0.030$ OOM
BONSAI GCN 0.3 $0.722\pm 0.011$ $0.585\pm 0.003$ $0.538\pm 0.103$ $0.683\pm 0.039$ $0.798\pm 0.007$
0.5 $0.701\pm 0.012$ $0.666\pm 0.001$ $0.670\pm 0.084$ $0.702\pm 0.049$ $0.832\pm 0.016$
GAT 0.3 $0.680\pm 0.014$ $0.566\pm 0.003$ $0.540\pm 0.096$ $0.665\pm 0.027$ $0.791\pm 0.010$
0.5 $0.683\pm 0.006$ $0.579\pm 0.004$ $0.523\pm 0.133$ $0.681\pm 0.025$ $0.812\pm 0.016$
FIT-GNN GCN 0.3 $0.800\pm 0.003$ $0.679\pm 0.003$ $0.754\pm 0.001$ $\mathbf{0.786\pm 0.001}$ $0.932\pm 0.000$
0.5 $\mathbf{0.829\pm 0.002}$ $0.668\pm 0.003$ $0.761\pm 0.004$ $0.700\pm 0.001$ $0.926\pm 0.000$
GAT 0.3 $0.761\pm 0.004$ $0.655\pm 0.005$ $0.754\pm 0.002$ $0.771\pm 0.002$ $0.885\pm 0.004$
0.5 $0.792\pm 0.003$ $0.669\pm 0.004$ $0.760\pm 0.003$ $0.720\pm 0.002$ $0.885\pm 0.003$
Node Regression: Table 3 shows the node regression error (mean and standard deviation of the lowest $10$ MAE). As the coarsening ratio increases, more subgraphs are created, and the size of each subgraph also reduces, and important neighboring node information gets lost. Hence, an increase in the coarsening ratio decreases performance.
Table 5: Results on OGBN-Products. FIT-GNN model uses a coarsening ratio $0.5$. variation_neighborhoods coarsening algorithm is used. OOM means Out Of Memory.
Full $0.873\pm 0.000$
SGGC OOM
GCOND OOM
BONSAI OOM
FIT-GNN $\mathbf{0.894\pm 0.000}$
Node Classification: Table 4 shows the results for node classification accuracy (mean and standard deviation of the top $10$ accuracies) with coarsening ratios $0.3$ and $0.5$ using the Cluster Nodes approach. Comparison with other coarsening ratios is in Table 10. Results show that the accuracy is comparable to the baselines. Figure 5 shows the comparison of performance between different experimental setups for different methods of appending nodes on the Cora dataset. From the Figure, we observe that Gc-train-to-Gs-train offers the best result among the three experimental setups for the node classification task. For the ’None’ method of appending nodes, where no additional nodes are added to each subgraph, we notice that it performs poorly compared to other methods, indicating that the Extra Node and Cluster Node methods indeed provide additional information for better empirical performance. We also observe in the ’None’ method that, with an increase in the coarsening ratio, the performance drops. This is because, as the coarsening ratio increases, the size of each subgraph gets smaller, hence more edges are cut due to partitioning, leading to more information loss. For the Gc-train-to-Gs-infer method, we observe that the performance drops as the coarsening ratio increases. This can be argued because of the change in the distribution of the coarsening graph and the set of subgraphs. When training on $G_{s}$, the distribution in the train and test sets is similar, hence attaining better performance. While comparing the methods of appending nodes, we infer that the Cluster Node method is performing similarly compared to the Extra Nodes method, if not better. Table 5 shows results on the OGBN-Products dataset. Result on training on the full dataset from the OGBN-Products leaderboard [Link]. For training on the full dataset, $8\%$ for training, next top $2\%$ for validation, and the rest for testing. For our approach, we use $0.038\%$ training, $0.057\%$ validation, and the rest for testing.
Table 6: Results on Graph Regression with MAE as metric (lower is better); We use Extra node and Gs-train-to-Gs-infer as experimental setup. variation_neighborhoods coarsening algorithm is used.
Methods Models Reduction Ratio (r) Dataset
ZINC (Subset) QM9 (prop 0) QM9 (prop 4) QM9 (prop 6) QM9 (prop 13)
Full GCN 1.0 $0.743$ $0.855$ $1.012$ $1.102$ $1.073$
GAT 1.0 $0.736$ $0.919$ $1.015$ $1.104$ $1.066$
SAGE 1.0 $0.685$ $0.885$ $1.015$ $1.109$ $1.076$
GIN 1.0 $0.748$ $0.925$ $1.053$ $1.126$ $1.078$
FIT-GNN GCN 0.1 $0.625$ $0.713$ $1.223$ $\mathbf{0.447}$ $1.223$
0.3 0.573 $0.708$ $0.975$ $0.473$ $1.220$
0.5 $0.645$ $0.702$ $1.153$ $0.485$ $1.179$
0.7 $0.651$ $0.715$ $1.189$ $0.456$ $1.240$
GAT 0.1 $0.574$ $0.705$ $0.940$ $0.464$ $0.524$
0.3 $0.652$ $0.708$ $1.056$ $0.528$ $0.564$
0.5 $0.620$ $0.693$ $1.227$ $0.475$ $0.561$
0.7 $0.658$ $\mathbf{0.686}$ $1.030$ $0.491$ $0.703$
SAGE 0.1 $0.735$ $0.726$ $\mathbf{0.739}$ $0.526$ $0.905$
0.3 $0.671$ $0.748$ $0.711$ $0.602$ $\mathbf{0.518}$
0.5 $0.695$ $0.739$ $0.747$ $0.606$ $0.833$
0.7 $0.645$ $0.722$ $0.677$ $0.659$ $1.005$
GIN 0.1 $0.644$ $0.726$ $0.746$ $0.494$ $0.856$
0.3 $0.669$ $0.723$ $0.761$ $0.509$ $0.854$
0.5 $0.631$ $0.725$ $0.848$ $0.737$ $1.058$
Graph Regression: Table 6 shows the results for the graph regression task on ZINC (Subset) and QM9 dataset. All the results show that the FIT-GNN model’s performance is better than the baselines. It is also observed that a lower coarsening ratio yields better loss, which implies that the model performs better on molecular graphs when the subgraph size is large. It implies that finer subgraphs lose out on global information about the molecule, which is necessary for the prediction. Similarly, when inferred on a coarsened graph for each graph in the dataset, the performance degrades. This is because, when a graph is reduced to a smaller graph, the positional information is lost, which is crucial since in graphs representing different molecules, the position of a certain component in a molecule results in different properties of the molecule.
Table 7: Results on Graph Classification with accuracy as metric (higher is better). We use the extra nodes method and Gc-train-to-Gc-infer as experimental setup. algebraic_JC coarsening algorithm is used.
Methods Models Graph per class Dataset
AIDS PROTEINS
DOSCOND GCN 1 $0.785$ $0.590$
10 $0.666$ $0.647$
50 $0.518$ $0.656$
GAT 1 $0.608$ $0.652$
10 $0.723$ $0.667$
50 $0.729$ $0.657$
KIDD GCN 1 $0.399$ $0.416$
10 $0.236$ $0.659$
50 $0.252$ $0.671$
GAT 1 $0.441$ $0.416$
10 $0.225$ $0.665$
50 $0.218$ $0.671$
SAGE 1 $0.653$ $0.416$
10 $0.361$ $0.650$
50 $0.283$ $0.673$
GIN 1 $0.759$ $0.412$
10 $0.637$ $0.655$
50 $0.634$ $0.674$
Methods Models
Reduction
Ratio (r)
Dataset
AIDS PROTEINS
Full GCN 1.0 $0.788$ $0.710$
GAT 1.0 $0.802$ $0.645$
SAGE 1.0 $0.770$ $0.613$
GIN 1.0 $0.800$ $0.774$
FIT-GNN GCN 0.1 $0.810$ $0.783$
0.3 $\mathbf{0.844}$ $\mathbf{0.826}$
0.5 $0.836$ $0.696$
0.7 $0.793$ $0.783$
GAT 0.1 $0.784$ $0.652$
0.3 $0.793$ $0.522$
0.5 $0.810$ $0.739$
0.7 $0.759$ $0.522$
SAGE 0.1 $0.793$ $0.696$
0.3 $0.828$ $0.696$
0.5 $0.793$ $0.696$
0.7 $0.819$ $\mathbf{0.826}$
GIN 0.1 $0.819$ $0.652$
0.3 $0.828$ $0.783$
0.5 $0.793$ $0.478$
0.7 $0.819$ $0.739$
Graph Classification: Table 7 shows graph classification results on AIDS and PROTEINS datasets, where training and inference are done on $G^{\prime}$. For baselines like DOSCOND and KIDD, we use ‘Graph per class’, whereas we use ‘Reduction Ratio (r)’ as a metric to denote how these algorithms reduce the size of training data. DOSCOND and KIDD propose to generate synthetic graphs that mimic the training data distribution. Therefore, to cover all the graphs in the dataset appropriately, this approach generates {1, 10, 50} graphs per class. However, the graph size is kept similar to the graph size in the training set. These approaches also don’t cover the test set. Compared to these baselines, our approach reduces each graph into a set of subgraphs and a coarsened graph based on a reduction ratio. As a result, each graph in the training set and test set gets reduced. From our results, we see that FIT-GNN outperforms all the baselines.
Table 11 and 12 present the ablation study on various coarsening algorithms for all the aforementioned tasks. From the results, we observe that variation_neighborhoods coarsening algorithm yields consistently better results compared to others. Therefore, we report results on all datasets using variation_neighborhoods coarsening algorithm.
6.2 Inference Time and Memory
As mentioned in Section 4.3, our inference time is less than the standard GNN if certain conditions are met. In Table 8(a) and 8(b), we club all baseline models into one, since all baselines are inferred on the whole graph. We empirically show the reduction of time and space by our FIT-GNN model during inference. For recording the inference time, we use the Python time package to calculate the difference in time before and after the inference step.
Table 8: Inference time (sec) (Lower is better) comparison for Node-Level tasks and Graph-Level tasks. We show single-node prediction time using two different coarsening ratios for Node-Level tasks. Here, we used Cluster Nodes for the node-level task. We used Gc-train-to-Gc-infer experimental setup for the graph-level tasks.
FIT-GNN
Dataset Baselines $\mathbf{r=0.1}$ $\mathbf{r=0.3}$
Chameleon $0.0027$ $0.0016$ $\mathbf{0.0014}$
Squirrel $0.0081$ $0.0017$ $\mathbf{0.0014}$
Crocodile $0.0070$ $0.0015$ $\mathbf{0.0015}$
Cora $0.0026$ $\mathbf{0.0019}$ $0.0020$
Citeseer $0.0031$ $\mathbf{0.0018}$ $0.0019$
Pubmed $0.0042$ $0.0019$ $\mathbf{0.0018}$
DBLP $0.0063$ $0.0020$ $\mathbf{0.0018}$
Physics Coauthor $0.0252$ $0.0020$ $\mathbf{0.0017}$
OGBN-Products $0.1762$ $0.0017$ $\mathbf{0.0016}$
(a) Node-Level Tasks
FIT-GNN
Dastaset Baselines $\mathbf{r=0.3}$ $\mathbf{r=0.5}$
ZINC (subset) $\mathbf{0.00184}$ $\mathbf{0.00184}$ $0.00190$
QM9 $\mathbf{0.00173}$ $0.00180$ $0.00191$
AIDS $0.00163$ $\mathbf{0.00155}$ $0.00163$
PROTEINS $0.00165$ $\mathbf{0.00160}$ $0.00163$
(b) Graph-Level Tasks
[…]
Figure 3: GPU memory consumption in MegaBytes (MB) (log scale) for FIT-GNN (Cluster Node) at different reduction ratios $r$ and Baseline, during inference. variation_neighborhoods coarsening algorithm is used.
In Table 8(a), we show the average time to predict for $1000$ queries (Prediction for Node $v_{i}$) in the Baseline and FIT-GNN models for node regression and classification tasks. The baseline model processes the entire graph, increasing inference time, especially for large graphs. In contrast, the FIT-GNN model only requires the relevant subgraph, resulting in faster predictions. For larger datasets like the subset of OGBN-Products, baseline inference is not feasible with our computational capacity. Therefore, to show a comparison, we take a subset of OGBN-Products with $165000$ nodes and $4340428$ edges. We see up to $\mathbf{100\times}$. Section C elaborates more on the inference time for datasets with a larger number of nodes. Traditional methods fail to infer for very large datasets due to memory constraints. However, our approach can give an inference on the whole graph. Figure 3 and Table 13 show the detailed comparison of memory consumption for different datasets for both Extra Nodes and Cluster Nodes methods, along with the baseline memory, which highlights that the FIT-GNN model uses up to $\mathbf{100\times}$ less memory than the Baseline. Figure 6 shows the time taken to coarsen for different coarsening ratios for the Cora dataset. As the coarsening ratio increases, the time also increases, since the number of subgraphs increases with the coarsening ratio.
Table 8(b) compares the inference time of graph classification and graph regression tasks. We predict for randomly selected $1000$ graphs from the test split. The table shows how our method is comparable and sometimes faster than the Baseline. Also, we observe that the inference time increases with a higher coarsening ratio. Because with a higher coarsening ratio, the number of nodes in $G^{\prime}$ increases, resulting in more edges. Overall, the inference time and memory for all the tasks mentioned are drastically less than the baselines while maintaining the performance.
7 Conclusion
In this paper, we have focused on inference time and memory and presented a new way to utilize existing graph coarsening algorithms for GNNs. We have provided theoretical insight corresponding to the number of nodes in the graph for which the FIT-GNN model reduces the time and space complexity. Empirically, we have shown that our method is comparable to the uncoarsened Baseline while being orders of magnitude faster in terms of inference time and consuming a fraction of the memory. A few possible future directions involve studying directed and weighted graphs, focusing on the theoretical connections between Extra Nodes and Cluster Nodes, and tackling real-world data-intensive challenges, for example, weather forecasting.
Appendix
Appendix A More on Extra Nodes and Cluster Nodes
A.1 Extra Nodes
Lemma 4.1. Models with $1$ layer of GNN cannot distinguish between $G$ and $\mathcal{G}_{s}$ when Extra Nodes method is used.
Proof.
There are $2$ sets of nodes in $G_{i}$ as follows:
  • The set $S_{1}$ of nodes with 1-hop neighbours in $G_{i}$.
  • The set $S_{2}$ of nodes with not all 1-hop neighbors in $G_{i}$
Let $\mathcal{I}^{1}_{i}$ be the number of nodes whose information does not get passed on after $1$ layer of GNN for $G_{i}$.
$\mathcal{I}^{1}_{i}=\left|\bigcup_{v\in S_{2}}\mathcal{N}_{1}(v)-V(G_{i})\right|$
[…]
$\Rightarrow\left|\mathcal{E}_{G_{i}}\right|=\mathcal{I}^{1}_{i}$
Hence, when Extra Nodes is used, 1 Layer GNN model cannot distinguish $G$ and $\mathcal{G}_{s}$. ∎
To understand the information loss when taking $2$ Layers of GNN, we divide $G_{i}$ into $3$ sets of nodes.
  • Nodes with 1-hop and 2-hop neighbours in $G_{i}$. Let us call this set $S_{1}$
  • Nodes with 1-hop neighbours in $G_{i}$ but $\exists v$ in 2-hop neighbourhood that is not in $G_{i}$. Let us call this set $S_{2}$
  • Node where $\exists v$ in 1-hop and 2-hop neighbourhood that is not in $G_{i}$. Let us call this set $S_{3}$.
Let $\mathcal{I}^{2}_{i}$ be the number of nodes whose information doesn’t get passed on after $2$ layer of GNN for $G_{i}$.
$\mathcal{I}^{2}_{i}=\left|\bigcup_{v\in S_{3}}\mathcal{N}_{2}(v)-V(G_{i})\right|$
When we use Extra Nodes, the information loss can be written as follows:
$\mathcal{I}^{2}_{i}=\left|\bigcup_{v\in S_{3}}\mathcal{N}_{2}(v)-V(G_{i})-\mathcal{E}_{G_{i}}\right|$
The above entity will depend on the density of the subgraphs formed and the number of connections each subgraph shares with each other. An algorithm with an objective to reduce this entity for all subgraphs will lose the least amount of information when Extra Nodes method is used.
A.2 Cluster Nodes
Given a partition matrix $P$, the features of the coarsened node are $X^{\prime}=P^{T}X$. Given a normalized partition matrix, the features of a node $v^{\prime}_{i}\in G^{\prime}$ is the degree-weighted average of the features of nodes in $C_{i}$. This is one of the functions $f$ used to create the features of the cluster node from $C_{i}$.
Previously, according to Lemma 4.1, there is no information loss when using the $1$ layer of GNN and Extra Nodes method. It was also easy to quantify in terms of the number of nodes. However, it is different for Cluster Nodes. Let us discuss the issues first.
  • Only a weighted version of node information is shared with the subgraph. Suppose $v_{c},v_{d}\in G_{i}$ is connected to $v_{a},v_{b}\in G_{j}$. Then the information contributed by these nodes is $\frac{d_{a}x_{a}+d_{b}x_{b}}{\sum_{p}d_{p}}$. Here $d_{p}$ represents the degree of node $v_{p}$, and $x_{p}$ represents the feature of node $v_{p}$.
  • Other node information will also be shared which is $\frac{\sum_{p\neq v_{a},v_{b}}d_{p}x_{p}}{\sum_{p}d_{p}}$. This will capture further dependencies.
The performance of Cluster Node will depend on some distance or similarity metric between $x_{c},x_{d}$ and $f(C_{j})$.
Appendix B More Algorithms
Algorithm 3 Training GNN (Train on $G^{\prime}$) (10.1145/3447548.3467256)
1:$G=(V,E,X)$; Labels $Y$; Node-Model $M$; Loss $\ell$; Number of Layers $L$
2:Apply a coarsening algorithm on $G$, and output a normalized partition matrix $P$;
3:Construct Coarsened graph $G^{\prime}$ using $P$;
4:Construct feature matrix for $G^{\prime}$ by $X^{\prime}=P^{T}X$;
5:Construct labels for $G^{\prime}$ by $Y^{\prime}=\arg\max(P^{T}Y)$ ;
[…]
8:Train $M$ to minimize LossLoss;
Algorithm 4 Node-Model(L,A,D,X)(L,A,D,X)
1:Number of Layers $L$; AA; $D$; $X^{(0)}=X$;
2:for $i=1$ to $L$ do
3:  $X^{(i)}=\sigma(\tilde{D}^{-\frac{1}{2}}\tilde{A}\tilde{D}^{-\frac{1}{2}}X^{(i-1)}\mathcal{W}^{(i-1)})$ $\quad\#$ Equation 1
4:end for
[…]
6:return $Z$
Algorithm 5 Graph-Model-$G^{\prime}(L,A^{\prime},D^{\prime},X^{\prime})$
1:Number of Layers $L;A^{\prime};D^{\prime};X^{(0)}=X^{\prime}$
2:for $i=1$ to $L$ do
[…]
4:end for
[…]
7:return $Z$
Appendix C More on Time and Space complexity
In Lemma 4.2, we provide the conditions that need to be satisfied for better space and time complexities.
For node-level tasks, to obtain lower FIT-GNN inference time complexity as compared to classical GNN, we want the following as per asymptotic analysis,
  1. 1.
    Baseline vs. Single-Node Inference:
    $\underset{\textup{\tiny Baseline Complexity}}{\underbrace{n^{2}d+nd^{2}}}\geq\underset{\textup{ \tiny FIT-GNN Complexity}}{\underbrace{\max_{i}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]}}$ (4)
  2. 2.
    Baseline vs. Full-Graph Inference:
    $\underset{\textup{\tiny Baseline Complexity}}{\underbrace{n^{2}d+nd^{2}}}\geq\underset{\textup{ \tiny FIT-GNN Complexity}}{\underbrace{\sum_{i=1}^{k}\bigl[(n_{i}+\phi_{i})^{2}d+(n_{i}+\phi_{i})d^{2}\bigr]}}$ (5)
Figure 4: The figure shows the feasibility of our inference methods for different node-level datasets against coarsening ratios. Two important inference setups for each dataset: single-node inference with the computational cost $\mathcal{O}\left(\max_{i}(\bar{n_{i}}^{2}d+\bar{n_{i}}d^{2})+n\right)$, and full-graph inference with computational cost $\mathcal{O}\left(\sum_{i=1}^{k}(\bar{n_{i}}^{2}d+\bar{n_{i}}d^{2})\right)$, are compared against the baseline (classical GNN) computational cost of $\mathcal{O}\left(n^{2}d+nd^{2}\right)$. The coarsening method used is variation_neighborhoods. Both axes are in logarithmic scale (base 10).
Figure 4 demonstrates the feasibility of our FIT-GNN method for different coarsening ratios with regards to full-graph inference time complexity. We empirically compute the LHS (baseline is classical GNN) and RHS of the Inequality 4 and 5for various coarsening ratios and plot the results for multiple datasets. The coarsening algorithm used is variation_neighborhoods.
We make the following inferences:
  1. 1.
    There is a clear dataset-dependent computational-cost trade-off across different coarsening ratios. Moreover, the minimum inference time for our approach occurs at different coarsening ratios for different datasets.
  2. 2.
    Under single-node inference setup, FIT-GNN demonstrates increasingly superior inference time performance as compared to classical GNN with increasing coarsening ratio. This is because the individual subgraph sizes decrease with increasing coarsening ratio.
  3. 3.
    We observe lower inference time complexity of Cluster_node method as compared to Extra_node method under both inference setups. This is expected as $|C_{G_{i}}|\leq|\mathcal{E}_{{G_{i}}}|,\ \forall i\in\{1,2,\dots,k\}$.
Appendix D Dataset Description
Dataset
Number of
Graphs
Average
Nodes
Average
Edges
Features Classes
PROTEINS $1113$ 1919 $72$ 33 22
AIDS $2000$ 77 $16$ $38$ 22
(a) Summary of datasets used for Graph classification Dataset
Dataset
Number of
Graphs
Average
Nodes
Average
Edges
Features
Number of
Targets
QM9 $130831$ 88 $18$ $11$ 1919
ZINC (subset) $10000$ $11$ $25$ 11 11
(b) Summary of datasets used for Graph regression Dataset
Dataset Nodes Edges Features Classes
Cora $2708$ $5278$ $1433$ 77
Citeseer $3327$ $4552$ $3703$ 66
Pubmed $19717$ $44324$ $500$ 33
DBLP $17716$ $52867$ $1639$ 44
Physics
Coauthor
$34493$ $247962$ $8415$ 55
OGBN-Products $2449029$ $61859140$ $100$ $47$
(c) Summary of datasets used for Node classification Dataset
Dataset Nodes Edges Features
Number of
Targets
Chameleon $2277$ $31396$ $128$ 11
Squirrel $5201$ $198423$ $128$ 11
Crocodile $11631$ $170845$ $128$ 11
(d) Summary of datasets used for Node regression Dataset
Appendix E Experiment Details (Parameters and Device Configuration)
For node classification and node regression tasks, we use Adam Optimizer with a learning rate of $0.01$ and $L_{2}$ regularization with $0.0005$ weight. We use Adam Optimizer with a learning rate of $0.0001$ and $L_{2}$ regularization with $0.0005$ weight for graph level tasks. For both coarsened graph (Gc-train) and set of subgraph based (Gs-train) model, we set epochs to 300 and the number of layers of GCN to 2 for training. We set the hidden dimensions for each layer of GCN to $512$.
The device configurations are Intel(R) Xeon(R) Gold 5120 CPU @ 2.202.20GHz, 256256GB RAM, NVIDIA A100 4040GB GPU. We use Pytorch Geometric to to train our models.
Appendix F Results
Figure 5 presents the ablation study conducted to compare the performance of FIT-GNN for different experimental setup using different methods for appending nodes in the subgraph. The experiments are conducted on the Cora dataset for different coarsening ratios.
Figure 5: The plot shows the ablation study conducted on the Cora dataset to determine which experimental setup performs better than the others. The plot also compares different methods of appending nodes to the subgraphs and how performance changes with varying coarsening ratios. variation_neighborhoods coarsening algorithm is used.
Figure 6 shows the time comparison for creating the subgraphs for Cora Dataset for different coarsening ratios. It is observed that the None method takes the least amount of time, which is obvious since the subgraphs are not appended with additional nodes. For the Extra Node Method, the time increases since the subgraph size increases as new nodes are appended. For Cluster Node, we observe the maximum time taken since not only are we looking at the neighboring nodes that are lost due to partition, but also we are keeping a track of the cluster to which these neighboring nodes belong, corresponding to which, we add the cross-cluster edges.
Figure 6: The plot shows how the coarsening time varies for the Cora dataset for different coarsening ratios using different methods of appending nodes to the subgraphs. variation_neighborhoods coarsening algorithm is used.
Table 10 presents the results on the node classification task with all the coarsening ratios $\{0.1,0.3,0.5,0.7\}$. To validate the use of “variation_neighborhoods” as a coarsening algorithm for all the results presented, we show an ablation study for different coarsening algorithms in Table 11 and 12. Table 13 shows the maximum GPU memory consumption during inference for different datasets with different coarsening ratios and methods of appending nodes. Note that this memory consumption is to store the graph and the weight matrix in the memory only.
Methods Models Reduction Ratio (r) Dataset
Cora Citeseer Pubmed DBLP Physics
Full GCN 1.0 $0.821\pm 0.002$ $0.706\pm 0.002$ $0.768\pm 0.020$ $0.724\pm 0.002$ $0.933\pm 0.000$
GAT 1.0 $0.809\pm 0.004$ $0.700\pm 0.003$ $0.740\pm 0.004$ $0.713\pm 0.004$ $0.909\pm 0.008$
SGGC GCN 0.1 $0.752\pm 0.002$ $0.704\pm 0.003$ $0.730\pm 0.004$ $0.728\pm 0.019$ $0.892\pm 0.010$
0.3 $0.808\pm 0.003$ $0.700\pm 0.001$ $0.773\pm 0.002$ $0.737\pm 0.011$ $0.928\pm 0.003$
0.5 $0.808\pm 0.001$ $0.716\pm 0.002$ $\mathbf{0.793\pm 0.002}$ $0.733\pm 0.008$ $0.931\pm 0.004$
0.7 $0.805\pm 0.002$ $0.711\pm 0.003$ $0.786\pm 0.003$ $0.733\pm 0.014$ $0.931\pm 0.003$
GAT 0.1 $0.630\pm 0.001$ $0.605\pm 0.025$ $89\pm 0.033 0.606$$06\pm 0.017$ $0.854\pm 0.008$
0.3 $0.686\pm 0.010$ $0.635\pm 0.021$ $0.773\pm 0.002$ $0.626\pm 0.028$ $0.870\pm 0.016$
0.5 $0.641\pm 0.026$ $0.635\pm 0.024$ $0.683\pm 0.024$ $0.604\pm 0.047$ $0.842\pm 0.018$
0.7 $0.643\pm 0.019$ $0.570\pm 0.041$ $0.733\pm 0.023$ $0.613\pm 0.018$ $0.851\pm 0.026$
GCOND GCN 0.1 $0.789\pm 0.003$ $0.676\pm 0.020$ $0.785\pm 0.003$ $0.718\pm 0.007$ $0.884\pm 0.008$
0.3 $0.806\pm 0.003$ $0.722\pm 0.001$ $0.779\pm 0.002$ $0.748\pm 0.001$ $0.876\pm 0.011$
0.5 $0.808\pm 0.004$ $\mathbf{0.730\pm 0.007}$ $0.773\pm 0.004$ $0.755\pm 0.002$ $0.804\pm 0.005$
0.7 $0.793\pm 0.000$ $0.680\pm 0.012$ $0.775\pm 0.002$ $0.751\pm 0.000$ $0.914\pm 0.004$
GAT 0.1 $0.657\pm 0.067$ $0.553\pm 0.116$ $0.511\pm 0.079$ $0.475\pm 0.057$ OOM
0.3 $0.549\pm 0.080$ $0.628\pm 0.095$ $0.370\pm 0.079$ $0.481\pm 0.034$ OOM
0.5 $0.712\pm 0.059$ $0.501\pm 0.085$ $0.409\pm 0.044$ $0.464\pm 0.030$ OOM
0.7 $0.706\pm 0.062$ $0.453\pm 0.118$ $0.416\pm 0.036$ $0.462\pm 0.033$ OOM
BONSAI GCN 0.1 $0.729\pm 0.008$ $0.576\pm 0.006$ $0.600\pm 0.098$ $0.699\pm 0.031$ $0.822\pm 0.014$
0.3 $0.722\pm 0.011$ $0.585\pm 0.003$ $0.538\pm 0.103$ $0.683\pm 0.039$ $0.798\pm 0.007$
0.5 $0.701\pm 0.012$ $0.666\pm 0.001$ $0.670\pm 0.084$ $0.702\pm 0.049$ $0.832\pm 0.016$
0.7 $0.705\pm 0.009$ $0.582\pm 0.007$ $0.474\pm 0.020$ $0.694\pm 0.023$ $0.844\pm 0.015$
GAT 0.1 $0.685\pm 0.004$ $0.589\pm 0.002$ $0.446\pm 0.014$ $0.642\pm 0.015$ $0.803\pm 0.005$
0.3 $0.680\pm 0.014$ $0.566\pm 0.003$ $0.540\pm 0.096$ $0.665\pm 0.027$ $0.791\pm 0.010$
0.5 $0.683\pm 0.006$ $0.579\pm 0.004$ $0.523\pm 0.133$ $0.681\pm 0.025$ $0.812\pm 0.016$
0.7 $0.703\pm 0.010$ $0.594\pm 0.002$ $0.497\pm 0.028$ $0.705\pm 0.047$ $0.799\pm 0.010$
FIT-GNN GCN 0.1 $0.790\pm 0.004$ $0.701\pm 0.001$ $0.757\pm 0.002$ $0.755\pm 0.001$ $0.909\pm 0.000$
0.3 $0.800\pm 0.003$ $0.679\pm 0.003$ $0.754\pm 0.001$ $\mathbf{0.786\pm 0.001}$ $\mathbf{0.932\pm 0.000}$
0.5 $\mathbf{0.829\pm 0.002}$ $0.668\pm 0.003$ $0.761\pm 0.004$ $0.700\pm 0.001$ $0.926\pm 0.000$
0.7 $0.813\pm 0.002$ $0.664\pm 0.001$ $0.756\pm 0.002$ $0.746\pm 0.001$ $0.925\pm 0.000$
GAT 0.1 $0.773\pm 0.021$ $0.694\pm 0.003$ $0.737\pm 0.002$ $0.729\pm 0.010$ $0.887\pm 0.003$
0.3 $0.761\pm 0.004$ $0.655\pm 0.005$ $0.754\pm 0.002$ $0.771\pm 0.002$ $0.885\pm 0.004$
0.5 $0.792\pm 0.003$ $0.669\pm 0.004$ $0.760\pm 0.003$ $0.720\pm 0.002$ $0.885\pm 0.003$
0.7 $0.797\pm 0.003$ $0.662\pm 0.003$ $0.759\pm 0.002$ $0.739\pm 0.002$ $0.903\pm 0.003$
Table 10: Results for node classification tasks with accuracy as the metric (higher is better). We use the Cluster Nodes method to append additional nodes to subgraphs and Gs-train-to-Gs-infer as experimental setup. variation_neighborhoods coarsening algorithm is used.
Cora Chameleon
Coarsening Method $\mathbf{r=0.1}$ $\mathbf{r=0.3}$ $\mathbf{r=0.1}$ $\mathbf{r=0.3}$
variation_neighborhoods $0.790\pm 0.004$ $0.800\pm 0.003$ $\mathbf{0.496\pm 0.005}$ $\mathbf{0.489\pm 0.003}$
algebraic_JC $\mathbf{0.827\pm 0.006}$ $\mathbf{0.804\pm 0.003}$ $0.544\pm 0.005$ $0.504\pm 0.003$
kron $0.758\pm 0.004$ $0.800\pm 0.002$ $0.571\pm 0.004$ $0.580\pm 0.013$
heavy_edge $0.736\pm 0.012$ $0.773\pm 0.006$ $0.572\pm 0.002$ $0.531\pm 0.002$
variation_edges $0.484\pm 0.006$ $0.471\pm 0.012$ $0.513\pm 0.003$ $0.542\pm 0.005$
variation_cliques $0.751\pm 0.012$ $0.801\pm 0.009$ $0.674\pm 0.009$ $0.679\pm 0.008$
Table 11: The table shows the FIT-GNN ablation study to compare various coarsening methods on Cora and Chameleon datasets, and the metrics for each dataset are accuracy (higher the better) and normalized MAE (lower the better), respectively.
PROTEINS ZINC (subset)
Coarsening Method $\mathbf{r=0.3}$ $\mathbf{r=0.5}$ $\mathbf{r=0.3}$ $\mathbf{r=0.5}$
variation_neighborhoods $0.652$ $\mathbf{0.739}$ $\mathbf{0.573}$ $\mathbf{0.620}$
algebraic_JC $\mathbf{0.826}$ $\mathbf{0.739}$ $0.823$ $0.810$
kron $0.522$ $0.652$ $0.663$ $1.201$
heavy_edge $0.565$ $0.609$ $2.114$ $1.227$
variation_edges $0.652$ $0.565$ $1.548$ $1.883$
variation_cliques $0.696$ $0.652$ $0.689$ $0.742$
Table 12: The table shows the FIT-GNN ablation study to compare various coarsening methods on PROTEINS and ZINC (subset) datasets. The metric for the PROTEINS dataset is accuracy (higher is better), and for the ZINC (subset) is normalized MAE (lower is better)
Dataset Appending Nodes FIT-GNN Baseline
$\mathbf{=0.1} r$$\mathbf{=0.3}$ $\mathbf{=0.5} r$$\mathbf{=0.7}$
Chameleon Cluster Node $0.201$ $0.235$ $0.277$ $0.418$ $2.078$
Extra Node $0.622$ $0.622$ $0.622$ $0.564$
Crocodile Cluster Node $1.127$ $1.132$ $1.654$ $1.779$ $10.937$
Extra Node $4.586$ $4.586$ $4.586$ $4.586$
Squirrel Cluster Node $1.436$ $1.532$ $1.815$ $2.869$ $8.614$
Extra Node $7.190$ $7.190$ $7.190$ $7.190$
Cora Cluster Node $1.249$ $0.618$ $0.661$ $0.827$ $14.992$
Extra Node $1.906$ $1.103$ $0.970$ $0.970$
Citeseer Cluster Node $30.097$ $2.048$ $1.195$ $1.607$ $47.170$
Extra Node $30.097$ $2.490$ $1.835$ $1.835$
Pubmed Cluster Node $0.584$ $0.542$ $0.544$ $0.546$ $39.166$
Extra Node $0.778$ $0.562$ $0.562$ $0.562$
DBLP Cluster Node $5.785$ $2.803$ $1.761$ $1.875$ $112.514$
Extra Node $28.271$ $6.008$ $2.308$ $2.207$
Physics Coauthor Cluster Node $22.911$ $12.481$ $12.767$ $13.688$ $1115.079$
Extra Node $41.066$ $19.936$ $19.936$ $19.936$
OGBN-Products Cluster Node $28.399$ $2840.706$
Table 13: Summary of maximum GPU memory consumption during inference of datasets used for node-level tasks. All units are in MegaBytes (MB)